### v1

# COBias and Debias: Minimizing Language Model Pairwise Accuracy Bias via Nonlinear Integer Programming

Abstract
For language model classification, would you prefer having only one workable class or having every class working? The latter makes more practical uses. Especially for large language models (LLMs), the fact that they achieve a fair overall accuracy by in-context learning (ICL) obscures a large difference in individual class accuracies. In this work, we uncover and tackle language models’ imbalance in per-class prediction accuracy by reconceptualizing it as the Contextual Oddity Bias (COBias), and we are the first to engage nonlinear integer programming (NIP) to debias it. Briefly, COBias refers to the difference in accuracy by a class A compared to its “odd” class, which holds the majority wrong predictions of class A. With the COBias metric, we reveal that LLMs of varied scales and families exhibit large per-class accuracy differences. Then we propose Debiasing as Nonlinear Integer Programming (DNIP) to correct ICL per-class probabilities for lower bias and higher overall accuracy. Our optimization objective is directly based on the evaluation scores by COBias and accuracy metrics, solved by simulated annealing. Evaluations on three LLMs across seven NLP classification tasks show that DNIP simultaneously achieves significant COBias reduction (-27%) and accuracy improvement (+12%) over the conventional ICL approach, suggesting that modeling pairwise class accuracy differences is a direction in pushing forward more accurate, more reliable LLM predictions.
1 Introduction
We rethink language models’ imbalance in per-class prediction accuracy and reconceptualize it as the Contextual Oddity Bias (COBias). In a nutshell, COBias refers to the difference in accuracy by a class A compared to its “odd” class, which holds the majority wrong predictions of class A. COBias reflects a common language model failure, seen in a variety of NLP benchmark tasks from text classification to question answering. Especially for large language models (LLMs), the fact that they achieve remarkable test accuracy over all class labels by few-shot in-context learning (ICL) [1] can obscure a large difference in individual class accuracies. It is a direct outcome of LLMs’ liking of specific patterns or answers, leading to highly imbalanced predicted class distributions [2, 3]. This per-class accuracy difference can be further attributed to ICL prompt sensitivity or rooted in the black-box pretraining data. To name a few, the causes encompass predilections for common tokens in the pretraining corpora [4], choices of prompt templates [5, 6, 7], demonstrations [8], and prompt orders [9, 10]. Unintended LLM behaviors may also exacerbate the bias, such as user-following sycophancy [11, 12, 13] and prompt toning effects [14, 15, 16, 17]. Despite the causes, COBias renders a language model less effective for lower-accuracy classes and hinders users from trusting the answers, heightening the need for a unified debiasing method at output level.
Importantly, it is not just that a model tends to over-predict a class, but more precisely, the model tends to predict certain classes (not always all others) as it. That is why we need to model the interaction between pairwise class predictions for better bias reduction. An example of an odd class and the per-class accuracy difference are shown by the upper plots in Figure 1. Essentially, mitigating per-class accuracy bias is a critical step in pushing forward LLM inference abilities, as it directly targets improving lower-accuracy classes while improving or at least not hurting the overall accuracy.
In this work, we first put forward COBias and propose the COBias evaluation metric, which measures the averaged pairwise class accuracy difference (Section 2). On three widely used, different-scale LLMs including GPT-2-XL [18], Llama-2-7B and Llama-2-13B [19], we manifest that large COBias consistently exists across seven varied NLP classification tasks.
To reduce COBias and rectify LLM outputs, we directly operate on the per-class probabilities given by the conventional ICL approach. We propose DNIP, the Debiasing as Nonlinear Integer Programming method, to adjust per-class probabilities towards lower COBias and higher overall accuracy, from a combinatorial optimization viewpoint (Section 3). We formulate the debiasing problem as a discrete correction weight selection problem. In particular, DNIP re-weights per-class probability with an optimal set of weights selected based on the evaluation metrics of accuracy and COBias, and thus obtains much less biased predictions with higher accuracy.
On three LLMs across seven NLP evaluation datasets, we show the effectiveness of DNIP. Our key key messages are as follows:
  • We introduce the COBias evaluation metric to assess the pairwise class accuracy bias in language model predictions, and reveal that LLMs of different scales and families consistently exhibit a large COBias score, at an average of 43%.
  • We propose Debiasing as Nonlinear Integer Programming (DNIP) to jointly minimize COBias and maximize accuracy. An efficient simulated annealing algorithm is adopted to solve the mathematical model.
  • Experiments show that DNIP achieves the best of both worlds for lower COBias (avg. $43\%\rightarrow 16\%$) and higher overall accuracy (avg. $52\%\rightarrow 64\%$) over the ICL approach.
2 The Contextual Oddity Bias and the Debiasing Problem
We conceptualize an odd class to be a class at inference time where predictions of another class are often biased towards it. An odd class is relative to the other class. The existence of odd classes manifests the per-class accuracy bias. We term the accuracy difference between a class A and its odd class as the Contextual Oddity Bias (COBias), suggesting a model’s tendency to over-predict an answer that is caught as odd than the correct answer. Notably, an odd answer may reflect surface patterns captured from the task-specific prompts or a common token in the pretraining data; therefore, even on the same task, which answers are odd can vary across different models. That’s why we use “contextual” to mean that the oddity of an answer depends on the model or task. In addition, the prefix “CO” in COBias can be interpreted as the pairwise nature of the bias, suggesting that if there is an odd class, there has to be a class that is biased to it.
The Contextual Oddity Bias Metric. We define the contextual oddity bias metric as follows. Formally, given an $N$-way classification task, let $x_{m}$ denote an input whose label is $y_{m}$. By few-shot prompting, let $\boldsymbol{\boldsymbol{p}_{m}}=(p_{m1},\dots,p_{mN})$ denote the output token probability over $N$ classes $c_{1},\dots,c_{N}$. The ICL prediction $\hat{y}_{m}$ is $\operatorname*{arg\,max}_{j\in\{1,\dots,N\}}\boldsymbol{p}_{m}$. For a given class $c_{i}$, we denote the class it is most biased towards as its oddest class $c_{o}$. Suppose test accuracy for the two is $A_{o}$ and $A_{i}$ respectively, then COBiassingle is computed by:
$\textrm{COBias}_{\textrm{single}}=\displaystyle\sum_{i\neq o}\bigg{|}A_{o}-A_{i}\bigg{|}$ (1)
Class accuracy $A_{i}$ is given by $\frac{1}{|\mathcal{C}_{i}|}\sum_{m\in\mathcal{C}_{i}}\mathbbm{1}\{\hat{y}_{m}=y_{m}\}$, where $\mathcal{C}_{i}$ is the set of indices of class $c_{i}$ examples, and $\mathbbm{1}(\boldsymbol{\cdot})$ returns 1 if the condition inside is satisfied and 0 otherwise.
However, bedsides the oddest class $c_{o}$, there could be the second oddest, the third oddest, or etc. class for a given class. Generalizing the definition, we extend the contextual oddity bias metric as:
$\textrm{COBias}={N\choose 2}^{-1}\displaystyle\sum_{i=1}^{N-1}\sum_{j=i+1}^{N}\bigg{|}A_{i}-A_{j}\bigg{|}$ (2)
The contextual oddity bias metric reflects the inter-class imbalance in accuracy. This metric considers each class as the odd class to others, representing the difference between every pair of classes.
The Debiasing Problem. The problem is two-fold - we need to reduce the contextual oddity bias, while maintaining or increasing the overall test accuracy.
3 Improving LLM Outputs by Debiasing as Nonlinear Integer Programming
Figure 1: An overview of the DNIP method. We take the TREC [20] task for illustration; the ICL approach uses 1-shot prompting whose prompt is given by light yellow text boxes. The upper-left plot shows that Loc. is an odd class to the others, holding their majority wrong predictions, especially for Ent.; the upper-right plot demonstrates the stark contrast in per-class accuracy, with 100% for Loc. and 0 for Ent.. The bottom plots show that DNIP greatly reduces pairwise class accuracy differences, especially boosting accuracy for Ent. ($0\rightarrow 66\%$).
In this work, we approach the debiasing problem as a nonlinear integer programming (NIP) problem. The objective is to optimize correction weights to rectify the biased, per-class probabilities, with directly modeling the COBias metric and two accuracy-related metrics. The weights are optimized on a set of labeled examples, and are simply plugged in to inference-time probabilities. An overview of the proposed DNIP method is shown in Figure 1. Our source code will be released.
3.1 The DNIP Objective
Considering $M$ examples with known class labels, we let $\boldsymbol{p}_{m}$ denote the normalized output probability over $N$ class tokens for an example, and $\boldsymbol{p}_{m}=(p_{m1},\dots,p_{mn},\dots,p_{mN})$, $\sum_{n=1}^{N}p_{mn}=1$, $m\in\{1,\dots,M\}$. To debias $\boldsymbol{p}_{m}$, we introduce a discrete $K$-point correction scale $\boldsymbol{\omega}$ ranging from 0 to 1, to re-weight class probability. Let $\boldsymbol{\omega}=(\omega_{1},\dots,\omega_{K})$ denote the parameter for the discrete weights in the correction scale, where $\omega_{k}=k/K$. An example $10$-point scale is given by $\boldsymbol{\omega}=(0.1,\dots,1.0)$. For each class probability in $\boldsymbol{p}_{m}$, it times a correction weight chosen from $\boldsymbol{\omega}$ to obtain a debiased class probability.
Finding optimal correction weights is a combinatorial optimization problem: for $N$-way class probability $\boldsymbol{p}_{m}$, each class selects a weight from the $K$-dimensional correction scale. To find the optimal combination, the brute force strategy is an enumerated search, where the number of weight combinations to explore is $K^{N}$ for an example; for $M$ examples, it results in a computational complexity of $\mathcal{O}(K^{N}\cdot M)$, which can not run in polynomial time, calling for smarter methods.
Instead, we model the debiasing problem as a nonlinear integer program, and search for an optimal set of class correction weights with simulated annealing. To this end, the problem boils down to selecting optimal weight indices, and we define an integer selection variable $\xi=(\xi_{1},\dots,\xi_{N})$ that selects a weight index $k$ for the $n$-th class probability in $\boldsymbol{p}_{m}$, namely:
$\xi_{n}=k\quad\text{if the $n$-th class selects }\omega_{k}.\$ (3)
The DNIP objective function and constraints are formally given by:
$\min\quad$ $Z(\xi)=\frac{1}{M}\displaystyle\sum_{m=1}^{M}\mathbbm{1}\{\hat{y}_{m}\neq y_{m}\}+\beta{N\choose 2}^{-1}\sum_{i=1}^{N-1}\sum_{j=i+1}^{N}\bigg{|}A_{i}-A_{j}\bigg{|}-\tau\displaystyle\sum_{j=1}^{N}\textrm{PMI}_{j}$ (4)
s.t. $\hat{y}_{m}=\operatorname*{arg\,max}_{n\in\{1,\dots,N\}}\omega_{\xi_{n}}p_{mn},\qquad m=1,\dots,M$ (5)
$A_{i}=\frac{1}{|\mathcal{C}_{i}|}\sum_{m\in\mathcal{C}_{i}}\mathbbm{1}\{\hat{y}_{m}=y_{m}\},\qquad i=1,\dots,N$ (6)
$\textrm{PMI}_{j}=\textrm{PMI}(\hat{S}^{j},S^{j})=\log\frac{f(\hat{S}^{j},S^{j})}{f(\hat{S}^{j})f(S^{j})},\qquad j=1,\dots,N$ (7)
$\xi_{n}\in\{1,\dots,K\},\qquad n=1,\dots,N$ (8)
where $\beta$ and $\tau$ are parameters that control the bias and the negative PMI terms respectively, $\hat{S}^{j}$ and $S^{j}$ denote the examples of prediction $j$ and true label $j$ respectively [^1] In details, $f(\hat{S}^{j})$ is the ratio between the number of examples with prediction $j$ and the total number of examples, similarly, $f({S}^{j})$ is the ratio between the number of examples with class label $j$ and the total number of examples, $f(\hat{S}^{j},S^{j})$ is the ratio between the number of correct predictions of class $j$ and the total number of examples. Therefore, $\textrm{PMI}_{j}=\log\big{(}\frac{\mathrm{n}(\hat{S}^{j},S^{j})/N}{\mathrm{n}(\hat{S}^{j})/N\cdot\mathrm{n}(S^{j})/N}\big{)}$, where $\mathrm{n}(\cdot)$ is the count. In actual computations, we apply add-$u$ smoothing to smooth out zeros in the numerator and denominator. The value of the smoothing coefficient $u$ is selected along other parameters on a development set.. The debiased class probabilities and prediction are:
$\boldsymbol{p}^{\ast}_{m}$ $=(\omega_{\xi_{1}}p_{m1},\dots,\omega_{\xi_{N}}p_{mN})$ (9)
$y^{\ast}_{m}$ $=\operatorname*{arg\,max}_{n\in\{1,\dots,N\}}\boldsymbol{p}^{\ast}_{m}$ (10)
The intuition is to encourage the adjusted predictions to have lower inter-class bias while improving overall accuracy. As the second term COBias has minimized the inter-class accuracy bias, the first term error rate brings the class prediction closer to the actual class, and the third term negative PMI further enforces the class prediction to be more confident when it is close to the actual class.
3.2 Solving DNIP with Simulated Annealing
Algorithm 1 Optimizing the selection of correction weights with simulated annealing
1:$(\xi,\boldsymbol{p},y,\omega)$: weight selection variable, class probabilities and labels, the weight scale
2:$(\xi,\{y^{\ast}_{m}\}_{1}^{M})$
3:$T\leftarrow T_{max}$, $\xi^{\ast}\leftarrow\xi\leftarrow$ INIT()
4:while $T\geq T_{min}$ do
5:     while inner-loop criterion is not satisfied do
6:         $\xi_{new}\leftarrow$ PERTURB($\xi$$\triangleright$ Sample a new $\xi$
7:         $\Delta z\leftarrow$ $z(\xi_{new})-z(\xi)$
8:         if $\Delta z\leq 0$ then
9:              $\xi\leftarrow$ $\xi_{new}$
10:              if $z(\xi_{new})<z(\xi^{\ast})$ then
11:                  $\xi^{\ast}\leftarrow$ $\xi_{new}$
12:         else if random($0,1$) $<$ exp($\Delta z/T$then
13:              $\xi\leftarrow$ $\xi_{new}$triangleright Accept a worse $\xi$
14:     $T\leftarrow$ $\alpha T$
15:$y^{\ast}\leftarrow$ INFER($\xi^{\ast},\boldsymbol{p},\omega$)
16:return $\xi^{\ast},y^{\ast}$
We solve DNIP by simulated annealing [21, 22]. Simulated annealing (SA) is a Metropolis-Hastings sampling algorithm, which is widely adopted to solve optimization problems with its versatility in dealing with highly nonlinear models, noisy data, and many constraints [23]. The DNIP mathematical model is such a case, pointing to the need of SA. We especially take advantages of SA algorithms’ searching ability - escaping from local optima and finding global optima after reaching local optima - to find optimal correction weights. In details, the SA strategy consists of searching the solution space starting from a randomly initialized $\xi$ and then generating a new one by perturbing it. The new solution is either accepted or rejected by the Metropolis criterion after evaluating the objectives.
Algorithm 1 describes the searching for optimal $\xi$ using SA. The outer loop performs the cooling/annealing process. In this work, a geometric decay with $\alpha=0.95$ and initial temperature of 200,000 is set as cooling schedule. The inner loop simulates the thermal equilibrium reaching process at a temperature, and its criterion ensures that either the number of generated or accepted solutions at a temperature is above a threshold [24]. We sample a new $\xi$ from the neighborhood by randomly substitute a selection in $\xi$. The computational complexity for the SA algorithm is $\mathcal{O}(NK)$. Please refer to the Appendix Section A for derivations.
4 Experiments
4.1 Experimental Setup
Evaluation Tasks. The proposed method is evaluated on a diverse set of binary and multi-class NLP classification tasks, across general domains and biomedical domains. The five general-domain evaluation datasets are 4-way news topic classification AGNews [25], 14-way ontology classification DBpedia [25], 5-way sentiment classification SST-5 [26], 6-way retrieval question classification TREC [27, 20], binary entailment recognition RTE [28]; the two domain-specific datasets from BLURB [29] are 5-way biomedical relation extraction DDI [30], and 3-way biomedical question answering PubMedQA [31]. The selected biomedical tasks are of practical uses. For example, detecting DDI is useful in preventing adverse effects from drug combinations.
Each dataset comes with a training and a test set. We make use of the default training set for DNIP optimization, except for AGNews, DBpedia, and DDI, for which we randomly select 10,000 training examples since the datasets are quite large. We then split the training examples into 95%, 5% optimization and development sets. We evaluate on the default test set except for AGNews, for which we evaluate on 5,000 randomly selected test examples. Evaluation metrics are COBias and accuracy.
Models and Experimental Configuration. We evaluate three open-source LLMs across varied scales and model families, including GPT-2-XL (1.5B parameters), Llama-2-7B (7B parameters), and Llama-2-13B (13B parameters). For ICL, we use 1-shot prompting, and obtain the output per-class softmax probabilities. To standardize calculations, we normalize the softmax probabilities over all classes. For DNIP, we tune parameters of $\beta,\tau,K,u$ on the development set. To account for variance caused by different demonstrations, for each model and dataset, we perform three runs of prompting with different 1-shot demonstrations and obtain three sets of initial probabilities. For both ICL and DNIP, we report the average test accuracy and COBias over the three runs. The demonstration is randomly selected from training examples. All prompting is done on an NVIDIA A100 GPU. The simulated annealing algorithm executes on CPU (execution time depends on the optimization set size, 3 to 30 minutes on AMD EPYC 7742 CPU or slightly longer on local laptop CPU).
Model Eval. Metric AGNews DBpedia SST-5 TREC RTE DDI PubMedaQA Avg.
1-shot ICL
GPT-2-XL
Acc
COBias
52.15.4
35.511.5
31.89.9
40.03.6
34.913.7
48.75.4
27.410.5
45.68.7
55.41.9
82.424.5
14.54.4
40.75.9
55.20.0
59.412.6
38.8
50.3
Llama-2-7B
Acc
COBias
86.42.5
14.06.5
88.92.0
13.52.1
42.111.1
55.61.5
66.76.6
33.210.0
66.34.3
61.610.5
6.70.4
41.41.7
40.36.7
40.916.1
56.8
37.2
Llama-2-13B
Acc
COBias
79.97.0
28.316.1
88.61.7
16.23.6
44.94.3
53.15.0
68.510.8
35.96.5
71.52.2
43.47.0
7.20.9
45.65.9
55.12.9
61.21.9
59.4
40.5
DNIP
GPT-2-XL
Acc
COBias
68.51.0
1.40.5
69.99.1
24.18.3
44.52.20
26.02.5
46.312.7
27.27.2
50.82.1
7.15.0
43.914.9
17.07.1
57.11.3
29.825.0
54.4
18.9
Llama-2-7B
Acc
COBias
86.70.4
1.30.1
92.90.4
7.70.6
50.62.7
28.021.6
68.11.0
15.91.6
73.92.3
1.91.8
44.53.8
11.63.0
62.78.3
35.422.8
68.5
14.5
Llama-2-13B
Acc
COBias
87.90.7
6.30.6
93.40.6
7.70.6
48.31.9
18.710.1
77.12.0
14.21.3
74.20.7
4.33.4
40.46.0
7.53.2
63.114.0
41.129.6
69.2
14.3
Table 1: Performance comparisons between the 1-shot ICL approach and DNIP. In each cell, we report the average test accuracy/COBias (%) and standard deviation over three different prompts. DNIP greatly improves test accuracy and significantly lowers test COBias upon the ICL approach.
4.2 Main Results
The average test accuracy (in black) and COBias (in blue) with standard deviation over three runs are shown in Table 1. We find that test COBias by ICL is large in all three LLMs, with GPT-2-XL having the largest average COBias of 50.3%, showing that LLMs of different sizes and families exhibit large per-class accuracy differences. With DNIP, the average test COBias over three models reduces from 43% to 16%, and the test accuracy increases from 52% to 64%. Specifically, DNIP achieves an 16%, 12%, and 10% absolute improvement in test accuracy for GPT-2-XL, Llama-2-7B, and Llama-2-13B respectively; meanwhile, DNIP significantly reduces test COBias by an absolute 29%, 23%, and 26% for the three models, over ICL results. As for detailed results, for GPT-2-XL where the ICL test accuracy is below 30% (DDI), DNIP boosts test accuracy by 29% while lowering test COBias from 41% to 17%; for Llama-2 7B and 13B where the ICL test accuracy is already in higher 80s, e.g., DBpedia, DNIP can further improve the accuracy to 90s while reducing COBias. In addition, we observe a very large ICL test COBias at 82.4% for GPT-2-XL on RTE. After applying DNIP, the COBias greatly reduces to 7.1%. The 82.4% COBias on RTE suggests that the conventional ICL predictions mostly fall in a single class, failing to predict the other class. In fact, The RTE task seems too challenging that GPT-2-XL’s output probability of class True is much higher than that of class False. Though re-weighting can lower both probabilities, it may not gain a corrected prediction. Still, the over 75% absolute reduction in COBias demonstrates that DNIP in the worst case can balance severely skewed ICL predictions without hurting the overall accuracy much. On two biomedical tasks, DNIP brings an average absolute increase of 34% and 11% in accuracy, and an average absolute reduction of 31% and 13% in COBias, showing its effectiveness for domain-specific tasks.
These results show that DNIP strongly mitigates COBias and greatly improves the accuracy of lower-performing classes, while boosting overall accuracy. Moreover, relatively large LLMs can suffer even higher COBias than smaller models on biomedical tasks, suggesting that the per-class accuracy bias does not go away as models scale, and debiasing is essentially needed to improve inference for larger LLMs.
4.3 DNIP Balances Class Accuracies across Varied Classification Tasks, as many as 14 Classes
Figure 2: Comparisons of pairwise class accuracy differences on test sets before and after applying DNIP; the pinker the higher difference, the greener the lower difference; average accuracy over three runs are used. Heatmaps after debiasing show nearly 0 differences for many class pairs, demonstrating DNIP’s effectiveness for COBias reduction.
We visualize pairwise class accuracy differences before and after applying DNIP on Llama-2-13B by the heatmaps in Figure 2, which show that DNIP is straightforwardly effective in reducing COBias. In each heatmap, the value of row $i$, column $j$ cell represents the absolute difference between the test accuracies of two classes $i,j$ , i.e., $|A_{i}-A_{j}|$. We observe that DNIP brings the accuracy differences down to close to zero for 4 out of 7 datasets, including AGNews, DDI, DBpedia, and RTE; and the most relative reduction is seen on TREC. These results show that DNIP consistently reduces prediction bias across different NLP classification tasks.
Figure 3: Test confusion matrix for DBpedia before and after applying DNIP. DNIP greatly reduces wrong predictions for Nat., which is the lowest-accuracy class before debiasing.
Relieving Odd Classes for 14-Way Ontology Classification. We further investigate the 14-way classification task DBpedia. Before debiasing, the Nat. class exhibits the largest class accuracy difference compared to other classes, suggesting that Nat. is either the highest-accuracy class manifesting a strong prediction bias towards Nat., or the lowest-accuracy class where predictions of true Nat. examples are biased to one or more odd classes. We plot the confusion matrix of test results of 1 run out of 3 runs in Figure 3 and find that Nat. has the lowest accuracy, and its predictions are biased to two odd classes, Trans. and Bldg., with Trans. being the oddest to Nat.. DNIP greatly reduces false negatives for Nat., i.e., reduces wrong Trans. and Bldg. predictions for Nat.. Besides these visualizations, on an average over three runs, the test accuracy for Nat. improves from 44% to 89%, and that for Trans. also increases from 78% to 87%, demonstrating DNIP’s effectiveness in reducing pairwise class accuracy bias for classification tasks of as many as 14 classes.
Furthermore, we see more balanced class accuracies with DNIP for smaller LLMs of Llama-2-7B and GPT-2-XL, in Appendix Section B.
4.4 DNIP Objective Ablation
We present ablation analysis on objective functions to show that the integration of accuracy-related objectives and the COBias objective are indispensable in achieving the best of both worlds. Recall three parts in our objective function $z$: the error rate term $z_{1}=\frac{1}{M}\sum_{m=1}^{M}\mathbbm{1}\{\hat{y}_{m}\neq y_{m}\}$; the COBias term $z_{2}={N\choose 2}^{-1}\sum_{i=1}^{N-1}\sum_{j=i+1}^{N}|A_{i}-A_{j}|$; and the PMI term $z_{3}=\sum_{j=1}^{N}\textrm{PMI}_{j}$; the accuracy-related terms are $z_{1}$ and $z_{3}$. We set up 7 different objective functions to perform DNIP: $z_{1},z_{2},z_{3},z_{1}+\beta z_{2},z_{1}-\tau z_{3},\beta z_{2}-\tau z_{3},z_{1}+\beta z_{2}-\tau z_{3}$, and compare their effectiveness on Llama-2-13B in Figure 4. Ablations on GPT-2-XL and Llama-2-7B show similar findings (Appendix Section C). To make fair comparisons, we fix $\beta=2.7,\tau=0.2$ and weight scale $K=30$, which may not result in the same scores reported in Table 1.
Figure 4: Ablation analysis of the objective function, exemplified with Llama-2-13B; test scores averaged over 7 datasets are used for the rightmost plot.
Among all ablations, $z_{1}+\beta z_{2}-\tau z_{3}$ achieves a balance point between accuracy and COBias. For the criteria, lower COBias is our top priority; when COBias scores are similarly low, we prefer the one that achieves higher accuracy. Compared to $z_{1}+\beta z_{2}-\tau z_{3}$$z_{1}$: (1) Using $z_{1}$ or $z_{3}$ solely emphasizes improving accuracy; though obtaining higher or similar accuracy, they suffer a higher COBias. (2) Using $z_{1}-\tau z_{3}$ also obtains a higher/similar accuracy but a higher COBias. (3) Using $z_{2}$ solely focuses on mitigating COBias and is straightforwardly effective in reducing COBias. However, the shortcoming is obvious: a lower accuracy. (4) Using $z_{1}+\beta z_{2}$ or $\beta z_{2}-\tau z_{3}$ is promising with a similar/lower accuracy and a slightly lower COBias. We note that accuracy-related objectives alone increase accuracy by over 16% ($75\%-59\%$), which is greater than using the objective combining COBias, suggesting the NIP can be a direction for those who look for accuracy improvement.
4.5 Performance Improvements vs. Number of Optimization Examples
We evaluate DNIP with varying optimization examples on Llama-2-13B. We find that DNIP requires only a small optimization set of 10 examples for a fair improvement in both COBias and accuracy, and it exhibits a further, emergent COBias reduction at several thousands optimization examples.
Figure 5: Relative test COBias reduction; higher is better.
Test COBias reduction reaches 9% to 72% at 10 optimization examples, and continues to grow with more optimization examples. Figure 5 shows the relative test COBias reduction over ICL results. At 10 examples, DNIP can achieve a relative COBias reduction ranging from 9% to 72%. We observe an overall trend of COBias reduction except for DDI and PubMedQA, who have a spike at 100 examples, but DDI rises again after 500 examples while PubMedQA levels off (PubMedQA full optimization size: 950). An explanation is that DNIP prioritizes COBias reduction more than accuracy increase at fewer examples; at more examples, the accuracy-related objectives show more influence. It also indicates that domain-specific datasets are more sensitive to the trade-off. We further notice a surge in reductions from 1,000 to full for DDI, SST-5, and DBpedia (full optimization sizes: 9,500, 8,116, 9,500), suggesting that around 8,000 examples are needed for a further emergent reduction - PubMedQA may see further COBias reduction if more optimization examples are available.
Figure 6: Relative test accuracy increase; higher is better.
Test accuracy is improved with only 10 optimization examples, and plateaus at around 500 examples. Figure 6 shows that, at 10 examples, the relative increase in accuracy ranges from 0.2% to 11% except DDI, which is 160%. At 100 examples, the increase ranges from 3% to 208% (DDI, $7.2\%\rightarrow 18.7\%$). The increase continues to grow and stabilizes at around 500 examples. At 1,000 examples, test accuracy for DDI rises to 52%, which is more than 6 times of the ICL accuracy, showing that DNIP is effective for this challenging biomedical NLP task. Moreover, we observe for RTE a declining trend of accuracy increase, which matches its growing trend of COBias reduction, reflecting the trade-off between the two objectives. We also note a slight decrease in accuracy improvement for SST-5 and DDI at full size, which matches their sudden increase in COBias reduction, verifying that DNIP has emergent COBias reduction abilities with thousands of optimization examples. Overall, DNIP improves both accuracy and COBias at 10 examples, and achieves acceptable COBias reduction and accuracy improvement at 100 examples, with test COBias reduced by 57% and test accuracy improved. For detailed scores, please refer to Appendix D.
5 Related Work
LLMs may fail in various ways. Analysis towards understanding LLM failure modes has gained remarkable insights on how they fail and how to identify/avoid failures. For example, understanding hallucinations in LLMs facilitates fact-checking evaluation metrics for LLMs.[32, 33, 34]. Jones and Steinhardt [14] reveals failure modes of LLMs which resemble cognitive biases and measures the failures on GPT-3.5 and Codex. Lu et al. [9] identify ICL order sensitivity that result in LLM prompting failures. Turpin et al. [10] shows that LLM reasoning abilities are affected by adding biasing features to the input. Along this line, we tackle a common LLM failure of the per-class accuracy difference, via nonlinear integer programming optimization on the evaluation metric scores.
6 Discussion
Using Chain-of-Thought Prompting to Improve LLM Predictions. Some may argue that a simple remedy for biased predictions is Chain-of-Thought (CoT) prompting, which elicit more reasoning in LLMs. However, Turpin et al. [10] points out that CoT explanations can be unfaithful, as they can be biased towards class A when prompts always put letter A as the answer, causing 36% accuracy drops on BBH [35]. We add that prompt-based models may fall for pattern matching instead of recalling learned factual knowledge [2]. Therefore, CoT may not help due to unfaithful explanations.
Using Calibration Methods to Mitigate ICL Biases. Previous calibration methods reduce biases caused by prompting. For example, Zhao et al. [4], Zhou et al. [36] use content-free or content-based prompts to measure an offset and calibrate class probabilities by removing it. This may be good for mitigating bias towards a single class, but it does not consider the interaction between class predictions, e.g., pairwise class accuracy bias studied in this work. Another hazard is using prompts to compute the offset, which may cascade biases brought by a certain or several prompt templates, and biases inherited from pretraining may not be reflected. Another line of works does not rely on further prompting. Han et al. [3] fits prototypical clusters on class probabilities and updates predictions by a cluster-label mapping. We share the same spirit of correcting biased predictions without more prompting, but we differ in that our correction weights are explicitly optimized on COBias and accuracy metrics, while the cluster-based decision boundary is implicitly learned.
Evaluating Closed LLMs. While closed LLMs such as ChatGPT and GPT-4 [37] are interesting to evaluate, they do not return logits/softmax probabilities over the entire vocabulary at an output token. For example, ChatGPT returns log probabilities of up to 20 most likely output tokens, and may not cover the probability for every class that is needed in our class prediction debiasing.
7 Conclusion
This work tackles the imbalance in per-class prediction accuracy of language models. We introduce COBias to assess pairwise class accuracy differences, and reveal that LLMs of varied scales and families exhibit large COBias at inference. To minimize COBias and push for higher accuracy, we view debiasing as a combinatorial optimization problem and propose DNIP, which optimizes class correction weights by jointly minimizing evaluation scores by COBias and maximizing accuracy. Empirical evaluations on three LLMs across seven multi-class NLP classification tasks demonstrate that DNIP simultaneously achieves lower test COBias and higher test accuracy over conventional ICL. We advocate that reducing COBias is a direction to gain more accurate, reliable LLM predictions.
Limitations and Future Work. The limitation is that we only evaluate NLP classification tasks. We leave evaluations on open-ended generations and other modalities for future work. Indeed, DNIP is fruitful for debiasing imbalanced class predictions, which can easily extend to vision/multimodal LMs, as long as there is a need for mitigating per-class accuracy differences. This simple inference-time plug-in solution is there for further boosting LM performances.
Appendix A Derivation of SA Algorithm Complexity
[…]
Appendix B Additional Heatmaps for Pairwise Accuracy Differences
Figure 7 and 8 demonstrate that more balanced class accuracies are achieved with DNIP for smaller LLMs of Llama-2-7B and GPT-2-XL.
Figure 7: Comparisons of pairwise test class accuracy differences before and after applying DNIP on GPT-2-XL; the pinker the higher accuracy difference, the greener the lower accuracy difference; average accuracy over three runs are used.
Figure 8: Comparisons of pairwise test class accuracy differences before and after applying DNIP on Llama-2-7B; the pinker the higher accuracy difference, the greener the lower accuracy difference; average accuracy over three runs are used.
Appendix C Additional Ablation Analysis
We show ablation analysis on GPT-2-XL and Llama-2-7B in Figure 9 and 10.
Figure 9: Ablations on objective function, GPT-2-XL, demonstrating that $z_{1}+\beta z_{2}-\tau z_{3}$ achieves a balance point between accuracy and COBias.
Figure 10: Ablations on objective function, Llama-2-7B, demonstrating that $z_{1}+\beta z_{2}-\tau z_{3}$ achieves a balance point between accuracy and COBias.
Appendix D Full Table with Varying Number of Optimization Examples for DNIP
The full results for all three models across varying optimization set sizes are shown in Table 2 (on the next page).
Opt. set Eval. Metric AGNews DBpedia SST-5 TREC RTE DDI PubMedaQA
GPT-2-XL
0
Acc
COBias
52.15.4
35.511.5
31.89.9
40.03.6
34.913.7
48.75.4
27.410.5
45.68.7
55.41.9
82.424.5
14.54.4
40.75.9
55.20.0
59.412.6
10$\ast$
Acc
COBias
72.82.1
17.21.7
40.65.2
28.06.6
40.33.4
40.11.8
51.55.7
39.71.0
50.30.8
6.57.6
23.51.0
40.63.5
53.70.6
41.80.0
100
Acc
COBias
70.05.0
14.81.3
41.429.9
27.88.9
44.31.3
30.83.4
49.010.1
31.15.9
49.82.5
5.34.9
42.25.7
17.17.1
46.710.8
29.216.6
500
Acc
COBias
68.21.2
2.71.8
44.533.2
25.79.6
43.41.5
29.83.7
49.98.5
31.46.2
50.52.4
5.23.9
45.69.2
24.514.3
49.911.4
17.715.8
1,000
Acc
COBias
69.61.0
3.01.2
42.032.2
27.611.0
42.82.3
28.25.9
50.99.0
24.55.4
50.32.2
5.23.3
48.34.9
22.92.5
same as
full
Full
Acc
COBias
68.51.0
1.40.5
69.99.1
24.18.3
44.52.20
26.02.5
46.312.7
27.27.2
50.82.1
7.15.0
43.914.9
17.07.1
57.11.3
29.825.0
Llama-2-7B
0
Acc
COBias
86.42.5
14.06.5
88.92.0
13.52.1
42.111.1
55.61.5
66.76.6
33.210.0
66.34.3
61.610.5
6.70.4
41.41.7
40.36.7
40.916.1
10$\ast$
Acc
COBias
86.42.5
14.06.5
89.91.4
12.52.1
51.40.9
36.23.6
70.10.9
22.45.4
74.92.1
4.85.1
31.721.0
26.78.2
44.50.6
28.73.9
100
Acc
COBias
88.40.4
5.80.6
91.80.7
9.71.0
50.91.5
34.312.7
70.11.0
16.72.4
73.62.6
2.90.7
44.92.5
21.06.6
62.64.5
35.418.1
500
Acc
COBias
86.80.9
2.80.3
92.10.6
8.61.6
50.81.7
35.812.3
69.71.2
15.22.1
74.32.9
3.11.4
69.31.7
34.50.3
63.57.5
37.320.4
1,000
Acc
COBias
86.80.3
1.90.5
92.50.2
8.00.4
51.02.1
30.920.4
69.50.8
15.71.2
74.02.5
2.61.2
56.57.9
29.53.4
same as
full
Full
Acc
COBias
86.70.4
1.30.1
92.90.4
7.70.6
50.62.7
28.021.6
68.11.0
15.91.6
73.92.3
1.91.8
44.53.8
11.63.0
62.78.3
35.422.8
Llama-2-13B
0
Acc
COBias
79.97.0
28.316.1
88.61.7
16.23.6
44.94.3
53.15.0
68.510.8
35.96.5
71.52.2
43.47.0
7.20.9
45.65.9
55.12.9
61.21.9
10$\ast$
Acc
COBias
86.01.9/
14.33.5
88.81.6/
14.72.5
49.20.8/
41.78.1
75.83.2
29.64.5
74.82.3
12.25.7
18.712.6
30.07.8
59.66.2
35.37.8
100
Acc
COBias
87.80.2/
8.30.5
91.80.2/
10.10.7
47.42.0/
30.44.4
78.32.8
18.52.6
75.51.0
10.46.9
22.22.0
17.34.4
56.910.4
25.816.3
500
Acc
COBias
87.11.5
5.33.1
92.70.0/
9.10.3
49.50.7/
28.03.2
78.72.1
16.31.0
73.61.7
4.03.4
48.227.4
28.011.5
62.215.8
40.429.4
1,000
Acc
COBias
87.70.7/
6.71.5
92.80.1/
9.20.2
49.80.7/
28.82.0
77.32.0
14.21.8
73.91.5
3.64.3
52.010.5
25.09.7
same as
full
Full
Acc
COBias
87.90.7
6.30.6
93.40.6
7.70.6
48.31.9
18.710.1
77.12.0
14.21.3
74.20.7
4.33.4
40.46.0
7.53.2
63.114.0
41.129.6
Table 2: Comparisons of DNIP results on varying optimization set sizes; min. optimization set size is 10 for all datasets except for DBpedia, which we use 15 to cover its 14 classes. For full optimization set sizes, AGNews: 9,500, DBpedia: 9,500, SST-5: 8,116, TREC: 5,179, RTE: 2,365, DDI: 9,500, PubMedQA: 950.