### v3

# Balancing Class Accuracies for Language Models in Inference Time via Nonlinear Integer Programming

Abstract
Large language models are good knowledge bases but struggle to perform equally well for all classes in simple text classification tasks. This paper investigates a fundamental inference-time problem in language models: imbalanced class accuracies. We find what’s underneath the issue is a tendency to over-predict some classes while under-predicting some others, for both base LLMs and more sophisticatedly trained, larger chat LLMs. This class accuracy imbalance is difficult to solve from the root via “better” pre-training or fine-tuning strategies, but we show it can be effectively mitigated via inference-time optimization. To this end, we conceptualize and quantify the over- and under-prediction issue as the Contextual Oddity Bias (COBias), and propose the Debiasing as Nonlinear Integer Programming (DNIP) model to correct in-context learned class probabilities based on minimizing COBias and maximizing overall accuracy, without LLM parameter update. Considering that the DNIP model implicitly contains non-differentiable elements, we therefore use the simulated annealing algorithm to solve. We perform extensive evaluations on three LLMs across seven NLP classification tasks, in different prompting settings. Results show that DNIP simultaneously achieves significant COBias reduction (-27%) and accuracy improvement (+12%) over the conventional ICL approach, suggesting that inference-time mitigation of class accuracy imbalance is a promising direction to push forward LLM performances.
1 Introduction
The fact that large language models (LLMs) achieve remarkable overall test accuracy by few-shot in-context learning (ICL) [1] can obscure a large difference in individual class accuracies. Mitigating class accuracy imbalance is under-explored, but important to facilitate a smoother user experience. Not mitigating such imbalances may lead to serious damages in some cases. For example, a patient can trust an LLM about drug-drug interaction type predictions without knowing that the model over-predicts “no interaction” and has low accuracies for other categories of interactions. If the patient trusts “no interaction” when the ground truth is an actual drug interaction type, taking the drugs together can put the patient’s life at risk. Therefore, it is high time we took actions to reduce class accuracy imbalances, better without hurting overall accuracy.
It is difficult to disentangle all the factors that cause class accuracy imbalance, including biases incurred in pre-training, supervised fine-tuning, RLHF, prompting specifics, etc. Therefore, it is challenging to comprehensively resolve the root causes through pre-training or fine-tuning methods, or approaches like data selection or data augmentation. To overcome these difficulties, we conceptualize a post-hoc bias metric, the Contextual Oddity Bias (COBias; Section 2), to quantify the class accuracy imbalance based on inference-time over-predictions and under-predictions.
To reduce COBias and rectify LLM outputs, we directly operate on the class probabilities obtained from in-context learning (ICL). We propose DNIP, the Debiasing as Nonlinear Integer Programming model, to adjust class probabilities towards lower COBias and higher overall accuracy, from a combinatorial optimization viewpoint (Section 3). We formulate the debiasing problem as a discrete correction weight selection problem. In particular, DNIP re-weights class probability with an optimal set of weights selected based on both metrics of accuracy and COBias, and thus obtains much less biased predictions with higher accuracy.
On three widely used different-scale LLMs across seven NLP evaluation datasets, and in different ICL settings, we show the effectiveness of DNIP. Our key messages are as follows:
  • We introduce the COBias evaluation metric to assess the pairwise class accuracy bias in language model predictions, and reveal that LLMs of different scales and families consistently exhibit a large COBias score.
  • We propose Debiasing as Nonlinear Integer Programming (DNIP) to jointly minimize COBias and maximize accuracy. An efficient simulated annealing algorithm is adopted to solve the mathematical model.
  • Experiments show that DNIP achieves the best of both worlds of lower COBias and higher overall accuracy. For example, on Llama-2-13B, DNIP significantly reduces COBias (avg. $40.5\%\rightarrow 14.3\%$) and improves overall accuracy (avg. $59.4\%\rightarrow 69.2\%$) over the 1-shot ICL results, and it reduces COBias even further over the 5-shot ICL results (avg. $39.2\%\rightarrow 10.3\%$).
2 Over-prediction, Under-prediction, and the Contextual Oddity Bias
2.1 Over-prediction and Under-prediction in Classification
[…]
In the era of Large Language Models (LLMs), over-prediction and under-prediction still persist in LLMs’ predictions, though one may not simply attribute such issues to the black-box training data or the model itself. The first obstacle is that the pretraining data’s source distribution for a subgroup of data may not be accessible, making it difficult to directly minimize differences between the source distribution and a downstream target distribution. What we normally have is the pretrained model’s output distribution for downstream classes. Secondly, the process of prompting an LLM, especially when the prompt is provided with in-context learned demonstrative examples, also brings uncertainty to predictions. Ablating the potential biases in prompting is non-trivial. More inconveniently, given the relatively large model sizes, it is computationally intensive, expensive, or even infeasible to tweak the LLM training pipeline or conduct LLM fine-tuning to fit debiasing training objectives.
To see how over-predictions and under-predictions hurt LLM prediction capabilities, we show in Table 2 an example of observed over-predictions and under-predictions in Llama-2-13B. We prompt the LLM on the AGNews news classification test set in a one-shot manner, with a single demonstrative example randomly selected from the training set. Llama-2-13B greatly over-predicts Business while under-predicting Technology, lowering the LLM’s capability for accurately predicting the Tech class (Tech’s accuracy: 19%). In general, there are 2129 predictions in Business but the number of ground-truth instances in Business is 1204; there are only 246 correct Tech predictions but 1240 ground-truth Tech instances. Over-/Under-prediction happens in more downstream tasks and larger LLMs (e.g., ChatGPT), as additionally show by examples in Appendix A. These surface observations tell us how an LLM could fail in predicting certain classes, suggesting that there is still space for improvements for LLMs to predict every class with equally good accuracy. The fact that this problem is difficult to completely eliminate from the source of data, together with the inconveniences in training/fine-tuning debiased LLMs, call for a more actionable inference-time bias metric quantifying the over-/under-predictions and an accompanying debiasing method.
Pred.
World Sports Business Tech All Acc.
World 1093 64 126 3 1286 0.85
Sports 9 1247 14 0 1270 0.98
True Business 25 4 1167 8 1204 0.97
Tech 156 27 822 235 1240 0.19
All 1283 1342 2129 246 5000 -
Prec. 0.85 0.93 0.55 0.96 - -
Table 1: Over-prediction in class Business and under-prediction in class Tech in Llama-2-13B, shown by the test confusion matrix for AGNews.
Pred.
World Sports Business Tech All Acc.
World 1100 50 67 69 1286 0.86
Sports 16 1240 11 3 1270 0.98
True Business 44 2 1028 130 1204 0.85
Tech 55 7 138 1040 1240 0.84
All 1215 1299 1244 1242 5000 -
Prec. 0.91 0.95 0.83 0.84 - -
Table 2: Significantly reduced over-predictions in class Business and improved test accuracy of the under-predicted class Tech for AGNews after DNIP.
2.2 Conceptualizing Over-prediction and Under-prediction into the Contextual Oddity Bias
In Table 2, we observe that Llama-2-13B greatly over-predicts Business while under-predicting Technology (Tech). Underneath the surface over-/under-prediction issue, we can see the most significant contributor is the 822 ground-truth instances of Technology, which are wrongly predicted as Business, leading to the model’s low accuracy (19%) in Technology [^1] It simultaneously leads to a fairly low precision score of 55% in Business. In fact, the low precision score in Business and the low recall/accuracy score in Technology co-reflect each other. In other words, class accuracies are essential indicators for class performances. All current evaluation metrics are computed by Python scikit-learn. Note that by scikit-learn, class recall is the same as class accuracy.. On that note, we are able to precisely describe the surface over-prediction and under-prediction issue, by decoupling the problem into which output class that a ground-truth class is most wrongly predicted as, suggesting the model’s inference bias towards an output class given a ground-truth class. In the AGNews example, Business is what ground-truth Technology is most wrongly predicted as, suggesting the model’s inference bias to predict Technology as Business.
To quantify this inference bias, we first conceptualize odd classes. An odd class is an output class at inference time which holds the most wrong predictions of a ground-truth class $c_{i}$. That is, an odd class is always relative to a ground-truth class. For $N$ classes and $M$ test instances, let $y$ and $\hat{y}$ denote the ground-truth class and the prediction respectively, and let $u_{c_{i},c_{j}}$ denote the number of ground-truth $c_{i}$ instances that are wrongly predicted as $c_{j}$,
$u_{c_{i},c_{j}}=\displaystyle\sum\nolimits_{m}\mathbbm{1}\{y_{m}=c_{i},\hat{y}_{m}=c_{j}\},$ (1)
$odd_class|(c_{i}):=\operatorname*{arg\,max}_{c_{j}}u_{c_{i},c_{j}},$ (2)
[…]
For this goal, it could be direct to use the rate of wrong predictions in an odd class as a measurement of its “hurt” to the accuracy of a ground-truth class, but minimizing these rates alone may not adequately mitigating class accuracy imbalances, which is an ultimate goal of reducing over-/under-predictions (More analyses on this measurement are in Appendix B.). A simpler yet better measurement is the absolute accuracy difference between the odd class and the ground-truth class. For example, knowing that Business is odd to Tech, the 78% accuracy difference between Business (97%) and Tech (19%) can serve as a measurement of how much Tech is under-predicted. The 19% accuracy of Tech also implicitly reflects the many predictions in its odd class Business. It should be noted that measuring these accuracy differences sometimes fails to capture the hidden ground-truth-and-odd relationships. For instance, World is the odd class to ground-truth Business, but the -12% accuracy difference can not utter this fact. Despite this drawback, directly minimizing the accuracy differences will in return significantly reduce predictions in odd classes and improve under-predicted classes’ accuracies.
The Contextual Oddity Bias. To this end, we formulate the definition of Contextual Oddity Bias (COBias). First, the absolute value of the accuracy difference between a ground-truth class and its odd class is termed as COBiassingle. COBiassingle measures a model’s tendency to over-predict an answer that is caught as odd than the correct answer (or equivalently, a tendency to under-predict the correct answer). Notably, for LLMs, an odd answer may be a manifest of surface patterns from prompts or biases in the pretraining data. In addition, the odd classes in a same downstream task can vary across different models. That’s why we use “contextual” to mean that the oddity of an answer depends on the model or task.
Formally, given an $N$ classes and $M$ instances, let $x_{m}$ denote an instance whose label is $y_{m}$. By few-shot prompting, the model’s output token probability over $N$ classes $c_{1},\dots,c_{N}$ is $\boldsymbol{p}_{m}=(p_{m1},\dots,p_{mN})$. $\boldsymbol{p}_{m}$ is normalized over $N$ classes, $\sum_{n=1}^{N}p_{mn}=1$, $m\in\{1,\dots,M\}$. Then the prediction $\hat{y}_{m}$ is $\operatorname*{arg\,max}_{j\in\{1,\dots,N\}}\boldsymbol{p}_{m}$. For a given class $c_{i}$, let $A_{\textrm{odd}_{i}}$ and $A_{c_{i}}$ denote the accuracies for the odd class $\textrm{odd}_{i}$ and $c_{i}$ respectively, the COBiassingle metric is computed as:
$\textrm{odd}_{i}=\verb|odd_class|(c_{i}),$ (3)
$\textrm{COBias}_{\textrm{single}}=\frac{1}{N}\displaystyle\sum\nolimits_{i}\bigg{|}A_{\textrm{odd}_{i}}-A_{c_{i}}\bigg{|},$ (4)
where class accuracy $A_{c_{i}}=\frac{1}{|\mathcal{S}_{c_{i}}|}\sum\nolimits_{m\in\mathcal{S}_{c_{i}}}\mathbbm{1}\{\hat{y}_{m}=y_{m}\}$, where $\mathcal{S}_{c_{i}}$ is the set of indices for class $c_{i}$ instances. Similarly, $A_{\textrm{odd}_{i}}=\frac{1}{|\mathcal{S}_{\textrm{odd}_{i}}|}\sum\nolimits_{m\in\mathcal{S}_{\textrm{odd}_{i}}}\mathbbm{1}\{\hat{y}_{m}=y_{m}\}$, where $\mathcal{S}_{\textrm{odd}_{i}}$ is the set of indices for class $\textrm{odd}_{i}$ instances.
Moreover, besides the odd class, which is the most-frequent wrongly predicted class, there could also be a secondly most-, or third most-, to the least-frequent wrongly predicted class. For a ground-truth class, there are the odd, the second odd, to the least odd classes. For example, for ground-truth Tech (1240 instances), Business is the odd (822), while World is the second odd (156), and Sports is the least odd (27). Then, we extend the accuracy difference between the single odd class and a ground-truth class to every pair of classes, and absolute value can be taken. This arrives at the ultimate the COBias metric as follows.
$\textrm{COBias}={N\choose 2}^{-1}\displaystyle\sum_{i=1}^{N-1}\sum_{j=i+1}^{N}\bigg{|}A_{c_{i}}-A_{c_{j}}\bigg{|}.$ (5)
This COBias metric directly reflects the inter-class imbalance in accuracy. Indeed, the absolute value of accuracy differences of some pairs in Equation 5 may mask the most frequent wrongly predicted class (the odd class). Nonetheless, it reflects a well-rounded measurement of all pairs of classes, whether involving odd classes or not, so it is a powerful metric when we strongly demand minimal class accuracy imbalance.
In a nutshell, the debiasing problem is minimizing COBias (Eq. 5) while maintaining or increasing the overall inference accuracy. This post-hoc bias evaluation metric will enable easy-to-operate inference-time/post-hoc output likelihood optimization, lifting the heavy computations required for training-wise data selection or model parameter updates.
3 Improving LLM Outputs by Debiasing as Nonlinear Integer Programming
Figure 1: An overview of the DNIP method. We take the TREC [5] task for illustration; the ICL approach uses 1-shot prompting whose prompt is given by light yellow text boxes. The upper-left plot shows that Loc. is an odd class to the others, holding their majority wrong predictions, especially for Ent.; the upper-right plot demonstrates the stark contrast in class accuracy, with 100% for Loc. and 0 for Ent.. The bottom plots show that DNIP greatly reduces pairwise class accuracy differences, especially boosting accuracy for Ent. ($0\rightarrow 66\%$).
In this work, we approach the debiasing problem as a nonlinear integer programming (NIP) problem. The objective is to optimize correction weights to rectify the biased, class probabilities, by directly minimizing the COBias metric while maximizing two accuracy-related metrics. The weights are optimized on a set of labeled examples, and are simply plugged in during test time. An overview of the proposed DNIP method is shown in Figure 1.
3.1 The DNIP Objective
To debias $\boldsymbol{p}_{m}$, we introduce a discrete $K$-point correction scale $\boldsymbol{\omega}$ ranging from 0 to 1, to re-weight class probability. Let $\boldsymbol{\omega}=(\omega_{1},\dots,\omega_{K})$ denote the parameter for the discrete weights in the correction scale, where $\omega_{k}=k/K$. An example $10$-point scale is given by $\boldsymbol{\omega}=(0.1,\dots,1.0)$. For each class probability in $\boldsymbol{p}_{m}$, it multiplies by a correction weight chosen from $\boldsymbol{\omega}$ to obtain a debiased class probability.
Finding optimal correction weights is a combinatorial optimization problem: for $N$-way class probability $\boldsymbol{p}_{m}$, each class selects a weight from the $K$-dimensional correction scale. To find the optimal combination, the brute force strategy is an enumerated search, where the number of weight combinations to explore is $K^{N}$ for an example; for $M$ examples, it results in a computational complexity of $\mathcal{O}(K^{N}\cdot M)$, which can not run in polynomial time, calling for smarter methods.
Instead, we model the debiasing problem as a nonlinear integer program, and search for an optimal set of class correction weights with simulated annealing. Without loss of generality, we boil down the problem to selecting optimal weight indices towards minimizing the pairwise class accuracy imbalance, while maximizing the overall accuracy and the pointwise mutual information between class labels and predictions. Formally, we define an integer selection variable $\xi=(\xi_{1},\dots,\xi_{N})$ that selects a weight index $k$ for the $n$-th class probability in $\boldsymbol{p}_{m}$, namely:
$\xi_{n}=k\quad\text{if the $n$-th class selects }\omega_{k}.\$ (6)
The DNIP objective function and constraints are formally given by:
$\min\quad$ $Z(\xi)=\frac{1}{M}\displaystyle\sum_{m=1}^{M}\mathbbm{1}\{\hat{y}_{m}\neq y_{m}\}+\beta{N\choose 2}^{-1}\sum_{i=1}^{N-1}\sum_{j=i+1}^{N}\bigg{|}A_{i}-A_{j}\bigg{|}-\tau\displaystyle\sum_{j=1}^{N}\textrm{PMI}_{j}$ (7)
s.t. $\hat{y}_{m}=\operatorname*{arg\,max}_{n\in\{1,\dots,N\}}\omega_{\xi_{n}}p_{mn},\qquad m=1,\dots,M$ (8)
$A_{i}=\frac{1}{|\mathcal{C}_{i}|}\sum_{m\in\mathcal{C}_{i}}\mathbbm{1}\{\hat{y}_{m}=y_{m}\},\qquad i=1,\dots,N$ (9)
$\textrm{PMI}_{j}=\textrm{PMI}(\hat{S}^{j},S^{j})=\log\frac{f(\hat{S}^{j},S^{j})}{f(\hat{S}^{j})f(S^{j})},\qquad j=1,\dots,N$ (10)
$\xi_{n}\in\{1,\dots,K\},\qquad n=1,\dots,N$ (11)
where $\beta$ and $\tau$ are parameters that control the bias and the negative PMI terms respectively, $\hat{S}^{j}$ and $S^{j}$ denote the examples of prediction $j$ and true label $j$ respectively. In details, $f(\hat{S}^{j})$ is the ratio between the number of examples with prediction $j$ and the total number of examples, similarly, $f({S}^{j})$ is the ratio between the number of examples with class label $j$ and the total number of examples, $f(\hat{S}^{j},S^{j})$ is the ratio between the number of correct predictions of class $j$ and the total number of examples. Therefore, $\textrm{PMI}_{j}=\log\big{(}\frac{\mathrm{n}(\hat{S}^{j},S^{j})/N}{\mathrm{n}(\hat{S}^{j})/N\cdot\mathrm{n}(S^{j})/N}\big{)}$, where $\mathrm{n}(\cdot)$ is the count. In actual computations, we apply add-$u$ smoothing to smooth out zeros in the numerator and denominator. The value of the smoothing coefficient $u$ is selected along other parameters on a development set.
The debiased class probabilities and prediction are:
$\boldsymbol{p}^{\ast}_{m}$ $=(\omega_{\xi_{1}}p_{m1},\dots,\omega_{\xi_{N}}p_{mN})$ (12)
$y^{\ast}_{m}$ $=\operatorname*{arg\,max}_{n\in\{1,\dots,N\}}\boldsymbol{p}^{\ast}_{m}$ (13)
The intuition is to encourage the adjusted predictions to have lower inter-class bias while improving overall accuracy. As the second term COBias has minimized the inter-class accuracy bias, the first term error rate brings the class prediction closer to the actual class, and the third term negative PMI further enforces the class prediction to be more confident when it is close to the actual class. The PMI term can also be seen as a constraint to penalize each individual class.
3.2 Solving DNIP with Simulated Annealing
Algorithm 1 Optimizing the selection of correction weights with simulated annealing
1:$(\xi,\boldsymbol{p},y,\omega)$: weight selection variable, class probabilities and labels, the weight scale
2:$(\xi,\{y^{\ast}_{m}\}_{1}^{M})$
3:$T\leftarrow T_{max}$, $\xi^{\ast}\leftarrow\xi\leftarrow$ INIT()
4:while $T\geq T_{min}$ do
5:     while inner-loop criterion is not satisfied do
6:         $\xi_{new}\leftarrow$ PERTURB($\xi$$\triangleright$ Sample a new $\xi$
7:         $\Delta z\leftarrow$ $z(\xi_{new})-z(\xi)$
8:         if $\Delta z\leq 0$ then
9:              $\xi\leftarrow$ $\xi_{new}$
10:              if $z(\xi_{new})<z(\xi^{\ast})$ then
11:                  $\xi^{\ast}\leftarrow$ $\xi_{new}$
12:         else if random($0,1$) $<$ exp($\Delta z/T$then
13:              $\xi\leftarrow$ $\xi_{new}$triangleright Accept a worse $\xi$
14:     $T\leftarrow$ $\alpha T$
15:$y^{\ast}\leftarrow$ INFER($\xi^{\ast},\boldsymbol{p},\omega$)
16:return $\xi^{\ast},y^{\ast}$
We solve DNIP by simulated annealing [6, 7]. Simulated annealing (SA) is a Metropolis-Hastings sampling algorithm, which is widely adopted to solve optimization problems with its versatility in dealing with highly nonlinear models, noisy data, and many constraints [8]. The DNIP mathematical model is such a case, pointing to the need of SA. We especially take advantages of SA algorithms’ searching ability - escaping from local optima and finding global optima after reaching local optima - to find optimal correction weights.
In details, the SA strategy consists of searching the solution space starting from a randomly initialized $\xi$ and then generating a new one by perturbing it. The new solution is either accepted or rejected by the Metropolis criterion after evaluating the objectives.
Algorithm 1 describes the searching for optimal $\xi$ using SA. The outer loop performs the cooling/annealing process. In this work, a geometric decay with $\alpha=0.95$ and initial temperature of 200,000 is set as cooling schedule. The inner loop simulates the thermal equilibrium reaching process at a temperature, and its criterion ensures that either the number of generated or accepted solutions at a temperature is above a threshold [9]. We sample a new $\xi$ from the neighborhood by randomly substitute a selection in $\xi$. The computational complexity for the SA algorithm is $\mathcal{O}(NK)$. Please refer to Appendix C for derivations.
4 Experiments
4.1 Experimental Setup
Evaluation Tasks. The proposed method is evaluated on a diverse set of binary and multi-class NLP classification tasks, across general domains and biomedical domains. The five general-domain evaluation datasets are 4-way news topic classification AGNews [10], 14-way ontology classification DBpedia [11], 5-way sentiment classification SST-5 [12], 6-way retrieval question classification TREC [13, 5], binary entailment recognition RTE [14]; the two domain-specific datasets from BLURB [15] are 5-way biomedical relation extraction DDI [16], and 3-way biomedical question answering PubMedQA [17]. The selected biomedical tasks are of practical uses. For example, detecting DDI is useful in preventing adverse effects from drug combinations.
Each dataset comes with a training and a test set. We make use of the default training set for DNIP optimization, except for AGNews, DBpedia, and DDI, for which we randomly select 10,000 training examples since the datasets are quite large. We then split the training examples into 95%, 5% optimization and development sets. We evaluate on the default test set except for AGNews, for which we evaluate on 5,000 randomly selected test examples. Evaluation metrics are COBias and accuracy.
Models and Experimental Configuration. We evaluate three open-source LLMs across varied scales and model families, including GPT-2-XL (1.5B parameters), Llama-2-7B (7B parameters), and Llama-2-13B (13B parameters). For main results, we use 1-shot prompting and obtain the ICL output class softmax probabilities. To standardize calculations, we normalize the softmax probabilities over all classes. For DNIP, we tune parameters of $\beta,\tau,K,u$ on the development set. To account for variance caused by different demonstrations, for each model and dataset, we perform three runs of prompting with different 1-shot demonstrations and obtain three sets of initial probabilities. For both ICL and DNIP, we report the average test accuracy and COBias over the three runs. The demonstration is randomly selected from training examples. All prompting is done on an NVIDIA A100 GPU. Simulated annealing executes on CPU (execution time depends on the optimization set size, 3 to 30 minutes on AMD EPYC 7742 CPU or slightly longer on local laptop CPU).
Model Eval. Metric AGNews DBpedia SST-5 TREC RTE DDI PubMedaQA Avg.
1-shot ICL
GPT-2-XL
Acc
COBias
52.15.4
35.511.5
31.89.9
40.03.6
34.913.7
48.75.4
27.410.5
45.68.7
55.41.9
82.424.5
14.54.4
40.75.9
55.20.0
59.412.6
38.8
50.3
Llama-2-7B
Acc
COBias
86.42.5
14.06.5
88.92.0
13.52.1
42.111.1
55.61.5
66.76.6
33.210.0
66.34.3
61.610.5
6.70.4
41.41.7
40.36.7
40.916.1
56.8
37.2
Llama-2-13B
Acc
COBias
79.97.0
28.316.1
88.61.7
16.23.7
44.94.3
53.15.0
68.510.8
35.96.5
71.52.2
43.47.0
7.20.9
45.65.9
55.12.9
61.21.9
59.4
40.5
DNIP
GPT-2-XL
Acc
COBias
68.51.0
1.40.5
69.99.1
24.18.3
44.52.20
26.02.5
46.312.7
27.27.2
50.82.1
7.15.0
43.914.9
17.07.1
57.11.3
29.825.0
54.4
18.9
Llama-2-7B
Acc
COBias
86.70.4
1.30.1
92.90.4
7.70.6
50.62.7
28.021.6
68.11.0
15.91.6
73.92.3
1.91.8
44.53.8
11.63.0
62.78.3
35.422.8
68.5
14.5
Llama-2-13B
Acc
COBias
87.90.7
6.30.6
93.40.6
7.70.6
48.31.9
18.710.1
77.12.0
14.21.3
74.30.8
4.33.3
40.46.0
7.53.2
63.114.0
41.129.6
69.2
14.3
Table 3: Performance comparisons between the 1-shot ICL approach and DNIP. In each cell, we report the average test accuracy/COBias (%) and standard deviation over three different prompts. DNIP greatly improves test accuracy and significantly lowers test COBias upon the ICL approach.
4.2 Main Results
The average test accuracy (in black) and COBias (in blue) with standard deviation over three runs are shown in Table 3. In all three LLMs, test COBias by ICL is large, with GPT-2-XL exhibiting the largest COBias (50.3%), showing that class accuracy imbalance exists in LLMs of different sizes and families. With DNIP, test COBias is significantly reduced, by an absolute 29%, 23%, and 26% for the three models respectively; meanwhile, there is an 16%, 12%, and 10% absolute improvement in test accuracy respectively. It is worth noting that for tasks whose ICL test accuracy is already in higher 80s, e.g., DBpedia, DNIP can reduce COBias while further improving the accuracy to 90s. In addition, we observe a very large ICL COBias at 82.4% in GPT-2-XL for RTE, reflecting severe over-predictions in class True. After applying DNIP, the COBias greatly reduces to 7.1%. That is, DNIP obtains over 75% absolute reduction in COBias for RTE, demonstrating DNIP’s capability for balancing very imbalanced class accuracies without hurting the overall accuracy much. Moreover, on the two biomedical tasks, DNIP brings an average absolute reduction of 31% and 13% in COBias, and an average absolute increase of 34% and 11% in accuracy, showing its effectiveness for domain-specific tasks.
Moreover, the 13B Llama model can suffer even higher COBias than its 7B counterpart on biomedical tasks, further showing that class accuracy imbalance does not go away as model sizes increase, and DNIP is effective to balance class accuracies for LLMs of all sizes.
Back to the AGNews example in Table 2, we present the DNIP debiased results in Table 2. Over-predictions in Business are greatly reduced, and the accuracy of Tech is significantly improved from 19% to 84%. These results show that DNIP strongly mitigates COBias and greatly improves the accuracy of lower-performing classes, while boosting overall accuracy.
Computational Time. Regarding computational time needed to perform DNIP optimization, on all datasets, the optimization is finished in the scale of minutes. Though computational time increases with more classes and a more fine-grained weight scale, it should be noted that the time elapsed in optimization is dozens of minutes even for tasks of more than 10 classes, and a weight scale of 30 is good enough for most of the datasets experimented in this paper. This demonstrates the practicality of DNIP in the computational time needed for performing optimization.
Prompting Method Metric AGNews DBpedia SST-5 TREC RTE DDI PubMedQA Avg.
Acc $82.5_{2.0}$ $93.6_{1.3}$ $45.8_{6.0}$ $58.7_{23.3}$ $61.9_{16.9}$ $34.1_{42.1}$ $70.4_{7.1}$ 63.9
5-shot ICL COBias $24.4_{4.2}$ $9.0_{2.0}$ $48.0_{15.4}$ $35.4_{13.7}$ $61.3_{40.4}$ $44.4_{5.4}$ $52.2_{19.3}$ 39.2
Acc $88.5_{0.6}$ $95.8_{0.7}$ $52.9_{2.5}$ $76.6_{8.7}$ $75.8_{4.5}$ $54.1_{5.1}$ $59.9_{1.4}$ 71.9
5-shot DNIP COBias $7.0_{0.9}$ $5.7_{1.1}$ $15.8_{12.6}$ $14.4_{7.5}$ $3.1_{1.9}$ $16.5_{7.6}$ $9.6_{4.6}$ 10.3
Acc $83.5_{1.5}$ $95.2_{1.2}$ $50.3_{2.3}$ $67.0_{12.7}$ $75.0_{0.8}$ $9.7_{1.0}$ $52.3_{5.3}$ 61.9
k-shot (1 demon. per class) ICL COBias $14.9_{5.1}$ $7.0_{2.2}$ $36.3_{7.2}$ $38.2_{5.1}$ $22.5_{13.2}$ $39.7_{3.5}$ $20.9_{4.2}$ 25.6
Acc $88.7_{0.5}$ $96.6_{0.5}$ $51.3_{1.0}$ $82.7_{1.4}$ $76.7_{3.7}$ $43.6_{6.1}$ $55.3_{2.6}$ 70.7
k-shot (1 demon. per class) DNIP COBias $7.3_{0.5}$ $4.3_{0.7}$ $2.8_{1.5}$ $12.1_{5.5}$ $5.0_{3.3}$ $12.5_{4.3}$ $8.7_{1.2}$ 7.5
Table 4: DNIP results in more ICL settings. Average score over three runs are reported.
4.3 More ICL Settings
We prompted Llama-2-13B in two additional settings: 5-shot, where each demonstration is randomly selected as in the 1-shot case; k-shot, where k is the number of classes and each class is represented by a demonstration example in the prompt. Table 4 presents results on seven benchmark datasets in both settings. For example, the relative average COBias reduction (comparing DNIP to ICL) is 65% and 74% for 1-shot and 5-shot cases respectively, and the relative average accuracy increase is 16% and 11% for 1-shot and 5-shot cases respectively, demonstrating DNIP’s capabilities with more shots. In addition, prompt design does contribute to different output probabilities, which could be helpful to provide different starting points for DNIP, so DNIP could optimize for better solutions. However, prompt engineering alone may not be most effective to solve the COBias issue. For example, compared to 1-shot prompting (Table 3), increasing the number of shots does not significantly further reduce COBias or boost accuracy; using more diverse demonstrations (k-shot) may help with COBias, but does not gain much higher accuracy than 1-shot. Therefore, using more shots does not help reduce COBias much, and DNIP significantly reduces COBias and improves accuracy in both 5-shot and k-shot settings, further showcasing its effectiveness.
4.4 DNIP Significantly Reduces Over-predictions, and Balances Class Accuracies across Varied Classification Tasks, as many as 14 Classes
Figure 2: Pairwise class accuracy differences on test sets before and after applying DNIP; the pinker the higher class accuracy difference, the greener the lower; average accuracy over three runs are used. Heatmaps after DNIP show nearly 0 accuracy differences for many class pairs.
Figure 3: Test confusion matrix for DBpedia before and after applying DNIP. DNIP greatly improves ICL predictions for Nat., which is the most under-predicted class by ICL.
We visualize pairwise class accuracy differences before and after applying DNIP to 1-shot, Llama-2-13B ICL results in Figure 2. In each heatmap, row $i$, column $j$ cell represents the absolute difference between the test accuracies of two classes in row $i,j$ , i.e., $|A_{i}-A_{j}|$. DNIP brings the accuracy differences down to close to zero for AGNews, DDI, DBpedia, and RTE. The lighter green heatmaps in the bottom row show that DNIP consistently reduces COBias across different NLP classification tasks.
Relieving Odd Classes for 14-Way Ontology Classification. We further investigate the 14-way classification task DBpedia. Before debiasing, the Nat. class exhibits the largest class accuracy difference to other classes. In fact, Nat. is the most under-predicted class, whose odd class is Trans. and second odd class is Bldg., as shown by the confusion matrix in Figure 3 (based on seed 0, 1-shot ICL run). DNIP greatly reduces false negatives for Nat., i.e., reduces wrong Trans. and Bldg. predictions for Nat.. Moreover, on an average over all three runs, the test accuracy for Nat. improves from 44% to 89%, and that for Trans. also increases from 78% to 87%, demonstrating that DNIP is effective in reducing class accuracy differences for tasks of as many as 14 classes. Furthermore, we present more heatmaps for smaller LLMs including Llama-2-7B and GPT-2-XL in Appendix D.
4.5 Optimizing COBiassingle
Using COBiassingle (Eq. 4) for DNIP optimization will focus on odd classes only. Figure 4 shows ablation results when replacing the COBias term in the DNIP objective with COBiassingle.
$\min\quad Z(\xi)=\frac{1}{M}\displaystyle\sum_{m=1}^{M}\mathbbm{1}\{\hat{y}_{m}\neq y_{m}\}+\beta\frac{1}{N}\displaystyle\sum\nolimits_{i}\bigg{|}A_{\textrm{odd}_{i}}-A_{c_{i}}\bigg{|}-\tau\displaystyle\sum_{j=1}^{N}\textrm{PMI}_{j}$ (14)
Figure 4: Test COBiassingle, COBias, and overall accuracy scores (from left to right).
Blue dots represent the respective evaluation metric score of the original 1-shot ICL Llama-2-13B outputs of the seven benchmark datasets; green dots are debiased results optimized by the COBiassingle DNIP objective (Eq. 14); red dots are debiased results optimized by the COBias DNIP objective.
Comparing DNIP with COBiassingle and DNIP with COBias, their performances are close, but DNIP with COBias achieves higher overall test accuracy and slightly better bias mitigation performances. In other words, DNIP with COBiassingle can significantly minimize accuracy differences between odd classes and ground-truth classes, but DNIP with COBias is even more effective when it comes to balancing all pairs of classes, not limited to the odd classes. In summary, optimizing either COBiassingle or COBias is an effective inference-time measure to reduce over-predictions in raw ICL outputs and boost accuracy for under-predicted classes.
4.6 Accuracy-related Objectives vs. the COBias Objective
We show that the integration of accuracy-related objectives and the COBias objective are indispensable in achieving the best of both worlds in Figure 5. DNIP objective function $z$ consists of three terms: the error rate term $z_{1}=\frac{1}{M}\sum_{m=1}^{M}\mathbbm{1}\{\hat{y}_{m}\neq y_{m}\}$; the COBias term $z_{2}={N\choose 2}^{-1}\sum_{i=1}^{N-1}\sum_{j=i+1}^{N}|A_{i}-A_{j}|$; and the PMI term $z_{3}=\sum_{j=1}^{N}\textrm{PMI}_{j}$; the accuracy-related terms are $z_{1}$ and $z_{3}$. We ablate all ablation objective functions to perform DNIP: $z_{1},z_{2},z_{3},z_{1}+\beta z_{2},z_{1}-\tau z_{3},\beta z_{2}-\tau z_{3},z_{1}+\beta z_{2}-\tau z_{3}$. Ablations on GPT-2-XL and Llama-2-7B show similar findings (Appendix E). To make fair comparisons, we fix $\beta=2.7,\tau=0.2$ and weight scale $K=30$, which may not result in the same scores reported in Table 3. Among all ablations, $z_{1}+\beta z_{2}-\tau z_{3}$ achieves a balance point between accuracy and COBias. For the criteria, lower COBias is our top priority; when COBias scores are similarly low, we prefer the one that achieves higher accuracy. Compared to $z_{1}+\beta z_{2}-\tau z_{3}$$z_{1}$: (1) Using $z_{1}$ or $z_{3}$ solely emphasizes improving accuracy; though obtaining higher or similar accuracy, they suffer a higher COBias. (2) Using $z_{1}-\tau z_{3}$ also obtains a higher/similar accuracy but a higher COBias. (3) Using $z_{2}$ solely focuses on mitigating COBias and is straightforwardly effective in reducing COBias. However, the shortcoming is obvious: a lower accuracy. (4) Using $z_{1}+\beta z_{2}$ or $\beta z_{2}-\tau z_{3}$ is promising with a similar/lower accuracy and a slightly lower COBias. We note that accuracy-related objectives alone increase accuracy by over 16% ($75\%-59\%$), which is greater than using the objective combining COBias, suggesting DNIP is effective for those who emphasize accuracy improvement.
Figure 5: Ablation analysis of the objective function, exemplified with Llama-2-13B; test scores averaged over 7 datasets are used for the rightmost plot.
4.7 Performance Improvements vs. Number of Optimization Examples
We evaluate DNIP with varying optimization examples on Llama-2-13B. We find that DNIP requires only a small optimization set of 10 examples for a fair improvement in both COBias and accuracy, and it exhibits a further, emergent COBias reduction at several thousands optimization examples.
Figure 6: Relative test COBias reduction; higher is better.
Test COBias reduction reaches 9% to 72% at 10 optimization examples, and continues to grow with more optimization examples. Figure 6 shows the relative test COBias reduction over ICL results. At 10 examples, DNIP can achieve a relative COBias reduction ranging from 9% to 72%. We observe an overall trend of COBias reduction except for DDI and PubMedQA, who have a spike at 100 examples, but DDI rises again after 500 examples while PubMedQA levels off (PubMedQA full optimization size: 950). An explanation is that DNIP prioritizes COBias reduction more than accuracy increase at fewer examples; at more examples, the accuracy-related objectives show more influence. It also indicates that domain-specific datasets are more sensitive to the trade-off. We further notice a surge in reductions from 1,000 to full for DDI, SST-5, and DBpedia (full optimization sizes: 9,500, 8,116, 9,500), suggesting that around 8,000 examples are needed for a further emergent reduction - PubMedQA may see further COBias reduction if more optimization examples are available.
Test accuracy is improved with only 10 optimization examples, and plateaus at around 500 examples. Figure 7 shows that, at 10 examples, the relative increase in accuracy ranges from 0.2% to 11% except DDI, which is 160%. At 100 examples, the increase ranges from 3% to 208% (DDI, $7.2\%\rightarrow 18.7\%$). The increase continues to grow and stabilizes at around 500 examples. At 1,000 examples, test accuracy for DDI rises to 52%, which is more than 6 times of the ICL accuracy, showing that DNIP is effective for this challenging biomedical NLP task.
Figure 7: Relative test accuracy increase; higher is better.
Moreover, we observe for RTE a declining trend of accuracy increase, which matches its growing trend of COBias reduction, reflecting the trade-off between the two objectives. We also note a slight decrease in accuracy improvement for SST-5 and DDI at full size, which matches their sudden increase in COBias reduction, verifying that DNIP has emergent COBias reduction abilities with thousands of optimization examples. Overall, DNIP improves both accuracy and COBias at 10 examples, and achieves acceptable COBias reduction and accuracy improvement at 100 examples, with test COBias reduced by 57% and test accuracy improved. For detailed scores, please refer to Appendix F.
5 Related Work
LLMs may fail in various ways. Analysis towards understanding LLM failure modes has gained remarkable insights on how they fail and how to identify/avoid failures. For example, understanding hallucinations in LLMs facilitates fact-checking evaluation metrics for LLMs.[18, 19, 20]. Jones and Steinhardt [21] reveals failure modes of LLMs which resemble cognitive biases and measures the failures on GPT-3.5 and Codex. Lu et al. [22] identify ICL order sensitivity that result in LLM prompting failures. Turpin et al. [23] shows that LLM reasoning abilities are affected by adding biasing features to the input. Along this line, we tackle a common LLM failure of the class accuracy difference, via nonlinear integer programming optimization on the evaluation metric scores. The inference-time class accuracy bias has multi-faceted causes. It could also be attributed to pre-training data, such as predilections for common tokens in the pretraining corpora [24]. It could be a direct outcome of LLMs’ liking for specific patterns or answers, leading to highly imbalanced predicted class distributions [25]. Other causes include choices of prompt templates [26, 27, 28], demonstrations [29], and prompt orders [22, 23]. Unintended LLM behaviors may also exacerbate the bias, such as user-following sycophancy [30, 31, 32] and prompt toning effects [21, 33, 34, 35].
6 Discussion
More Methods That Potentially Improve Accuracy while Reducing COBias. Prior methods tackle the prediction bias problem from different angles. Although COBias is not be explicitly modeled, they are promising to reduce COBias. Therefore, we draw comparisons to two popular categories of debiasing methods, adaptations and calibrations. In particular, with the same class probability vectors as input, we build a single linear adapter (LA) network with adapter dimension of 64, and train it towards the same objective as DNIP; we also follow the batch calibration method [36] to estimate a calibration term using test examples.
Figure 8: Comparisons to more methods.
Figure 8 shows the comparison results. fine-tuning a simple linear adapter does not work well on the class output probabilities. The proposed method outperforms both methods in accuracy and COBias.
Using Chain-of-Thought Prompting to Improve LLM Predictions. Some may argue that a simple remedy for biased predictions is Chain-of-Thought (CoT) prompting, which elicit more reasoning in LLMs. However, Turpin et al. [23] points out that CoT explanations can be unfaithful, as they can be biased towards class A when prompts always put letter A as the answer, causing 36% accuracy drops on BBH [37]. We add that prompt-based models may fall for pattern matching instead of recalling learned factual knowledge [25]. Therefore, CoT may not help due to unfaithful explanations.
Using Calibration Methods to Mitigate ICL Biases. Previous calibration methods reduce biases caused by prompting. For example, Zhao et al. [24], Zhou et al. [36] use content-free or content-based prompts to measure an offset and calibrate class probabilities by removing it. This may be good for mitigating bias towards a single class, but it does not consider the interaction between class predictions, e.g., pairwise class accuracy bias studied in this work. Another hazard is using prompts to compute the offset, which may cascade biases brought by a certain or several prompt templates, and biases inherited from pretraining may not be reflected. Instead, we correcct biased predictions without more prompting, but we differ in that our correction weights are explicitly optimized on COBias and accuracy metrics, while the cluster-based decision boundary is implicitly learned.
Evaluating Closed LLMs. While closed LLMs such as ChatGPT and GPT-4 [38] are interesting to evaluate, they do not return logits/softmax probabilities over the entire vocabulary at an output token. For example, ChatGPT returns log probabilities of up to 20 most likely output tokens, and may not cover the probability for every class that is needed in our class prediction debiasing.
Optimization Algorithms for Integer Programming. The classic solvers for integer programming include operational research techniques such as Branch and Bound (BnB), commonly used for linear integer programming problems. However, it is difficult for these methods to solve our DNIP model, which contains discontinuous functions and is non-differentiable. In this case, a series of metaheuristic algorithms are more suitable, solversincluding Simulated Annealing (SA), Genetic Algorithms, Ant Colony Optimization, and Particle Swarm Optimization. These algorithms all belong to the same category, and each meta heuristic has their own pros and cons. Therefore, SA in this paper has a clear purpose of tackling our NIP mathematical model that classical operational research methods cannot solve.
Our Motivation Is Different from Classical Post-hoc Corrections. Some may contend that achieving equitable accuracies across all classes is a well-explored issue in standard machine learning classifiers. However, it is crucial to see that class prediction accuracy imbalances must be understood within their specific context. The accuracy bias in LLMs’ outputs arises from fundamentally different causes than the unequal class accuracies seen in potentially overfitted traditional classifiers. In the former, biases stem from prompts and pre-training, while in the latter, the imbalance is driven by skewness in the supervised training data. This distinction underscores why our approach is uniquely applied to LLMs’ output token class probabilities.
7 Conclusion
This work tackles the imbalance in class prediction accuracy of language models. We introduce COBias and reveal that LLMs of varied scales and families exhibit large COBias at inference. To minimize COBias and push for higher accuracy, we view debiasing as an inference-time combinatorial optimization problem and propose DNIP, which optimizes class correction weights by jointly minimizing evaluation scores by COBias and maximizing accuracy. Empirical evaluations on three LLMs across seven multi-class NLP classification tasks demonstrate that DNIP simultaneously achieves lower test COBias and higher test accuracy over conventional ICL. We advocate that reducing COBias is a direction to gain more accurate, reliable LLM predictions. For future works, we may extend our studies to more tasks and other modalities. We aim to find simple inference-time plug-in solution for further boosting language model prediction performances.
Appendix A Other Examples of Over-prediction and Under-prediction in LLMs
We present Llama-2-13B’s and ChatGPT’s over- and under-predictions on a class imbalanced classification task, the Drug-Drug Interaction (DDI) relation classification task, where around 90% test instances are Negative. As shown in Table 5, Llama-2-13B significantly under-predicts Negative. In Table 6, similar under-predictions in the DDI Negative class happens for ChatGPT (GPT-3.5-turbo-instruct [39]), the RLHF’d chat LLM with magnitudes of more parameters. (It is noted that at the time of our experiments, ChatGPT does not support returning full probabilities over the entire vocabulary. To investigate the over-/under-prediction issue, we prompt ChatGPT on the DDI task with the following prompt template: “Please choose one most suitable answer for the following description. [ONE-SHOT DEMONSTRATION] Description: [TEST INPUT] Q: What is the drug-drug interaction relation between the aforementioned @DRUG\$ pair? Choose from Negative, Mechanism, Effect, Advice, Interaction. A: ”, and use the answer token directly as the prediction, and plot the confusion matrix in Table 6. For fair comparisons, we prompt Llama-2-13B in the same manner to obtain Table 5, which is different from how we prompt Llama-2-13B in the main paper.)
Pred.
Neg Eff Mech Adv Int
Neg 1749 438 1498 546 506
Eff 118 33 110 52 47
True Mech 77 28 123 45 29
Adv 84 15 62 36 24
Int 34 11 28 13 10
Table 5: Over-prediction and under-prediction in Llama-2-13B, as shown by the confusion matrix on DDI test set.
Pred.
Neg Eff Mech Adv Int
Neg 809 707 1001 217 2003
Eff 45 109 79 6 121
True Mech 15 46 172 0 69
Adv 24 12 18 79 88
Int 9 9 7 1 70
Table 6: Over-prediction and under-prediction in ChatGPT, as shown by the confusion matrix on DDI test set.
The over-/under-predictions in LLMs of different sizes and families, and the inconveniences in training/fine-tuning LLMs, together call for a more actionable bias metric quantifying the over-/under-prediction issue and an accompanying debiasing method tailored for LLMs.
Appendix B Directly Optimizing the Odd Prediction Rate
We discuss another option of measuring class accuracy imbalances, which does not mitigate odd predictions as effective as using COBias, though using it can obtain similar or slightly higher overall accuracy. It is the odd prediction rate, i.e., the rate of odd predictions over the number of ground-truth instances, which is a measurement reflecting the odd class and the level of “accuracy hurt” caused by the odd class. We define $H(c_{i})$ as below.
$\textrm{odd}_{i}=\verb|odd_class|(c_{i}),$ (15)
$H(c_{i})=\frac{\displaystyle\sum\nolimits_{m}\mathbbm{1}\{y_{m}=c_{i},\hat{y}_{m}=\textrm{odd}_{i}\}}{\displaystyle\sum\nolimits_{k}\mathbbm{1}\{y_{k}=c_{i}\}},$ (16)
where $\textrm{odd}_{i}$ is the odd class to $c_{i}$. For example, one may quantify the effect of a current odd class Business to the ground-truth class Tech as the rate of Business predictions over ground-truth Tech instances, i.e., 822/1240, and an immediate debiasing goal is to minimize this rate. The DNIP objective with the odd prediction rate is as follows.
$\min\quad Z(\xi)=\frac{1}{M}\displaystyle\sum_{m=1}^{M}\mathbbm{1}\{\hat{y}_{m}\neq y_{m}\}+\beta\frac{1}{N}\displaystyle\sum\nolimits_{i}H(c_{i})-\tau\displaystyle\sum_{j=1}^{N}\textrm{PMI}_{j}$ (17)
We compare the effectiveness of using COBias, COBiassingle, and $H$ in the DNIP objective, and refer to the respective objectives as ZCB (Eq. 4), ZCBs (Eq. 14), and ZH (Eq. 17). Detailed test scores for the benchmark datasets on Llama-2-13B, including class accuracies, are shown in Tables 7 to 13.
In short, ZH still leaves some classes under-predicted, as examplified by the individual test class accuracies, and test COBias (or test COBiassingle) scores on TREC, RTE, etc. The rationale is that optimizing bias measurements that extend accuracy differences, such as the odd prediction rates, does not mitigate odd class predictions better than optimizing COBias directly. ZCB most effectively mitigates the over- and under-prediction issue, balancing class accuracies.
Optimizing other extended measurements, such as precision or $F_{1}$ scores differences, may obtain similar performances to the three existing metrics, which will be left for future discussions.
DNIP Objective (Results without DNIP are shown on the last row.) Evaluation Metric
CB CBs H Overall Acc. Class Acc.
Co. Sch. Art. Ath. Polit. Trans. Bldg. Nat. Vill. Anim. Plt. Alb. Film Bk.
ZCB 7.7 7.5 3.3 93.4 82.6 97.7 94.0 99.5 97.9 94.4 86.6 89.0 99.5 96.3 98.5 98.6 97.0 77.8
ZCBs 9.1 5.7 3.9 92.6 87.4 96.8 97.1 98.9 98.1 87.3 87.6 84.0 99.3 99.2 97.9 99.6 96.5 69.0
ZH 8.9 9.5 2.8 93.2 76.2 99.3 95.0 99.4 98.8 96.9 82.8 89.8 99.5 98.4 98.9 99.8 96.1 76.3
ICL 16.2 13.7 5.6 88.6 91.7 99.1 92.7 99.7 97.0 94.8 78.4 44.2 99.7 96.9 99.3 99.8 95.2 63.3
Table 7: Comparison of DBpedia test set evaluation scores (%) for three DNIP objective functions ZCB, ZCBs, and ZH. Average scores over three runs of different 1-shot demonstrations are reported.
DNIP Objective (Results without DNIP are shown on the last row.) Evaluation Metric
CB CBs H Overall Acc. Class Acc.
Wrd. Sports Bus. Tech.
ZCB 6.3 4.4 7.2 87.9 85.3 95.9 86.5 83.9
ZCBs 7.3 4.3 6.9 88.4 84.8 98.2 85.8 84.7
ZH 8.5 4.8 6.5 88.3 86.8 98.7 84.2 83.2
ICL 28.3 22.5 14.6 79.9 81.5 98.7 91.0 48.0
Table 8: Comparison of AGNews test set evaluation scores (%) for three DNIP objective functions ZCB, ZCBs, and ZH. Average scores over three runs of different 1-shot demonstrations are reported.
DNIP Objective (Results without DNIP are shown in the last row.) Evaluation Metric
CB CBs H Overall Acc. Class Acc.
Neg. Eff. Mech. Adv. Int.
ZCB 7.5 7.0 39.6 40.4 42.1 31.6 32.8 33.9 31.6
ZCBs 19.9 16.8 42.1 22.9 19.9 34.2 37.3 39.8 44.8
ZH 43.6 44.6 30.9 21.7 17.6 12.6 47.2 85.5 33.3
ICL 45.6 71.8 59.4 7.2 0.0 30.7 61.5 24.1 43.5
Table 9: Comparison of DDI test set evaluation scores (%) for three DNIP objective functions ZCB, ZCBs, and ZH. Average scores over three runs of different 1-shot demonstrations are reported.
DNIP Objective (Results without DNIP are shown in the last row.) Evaluation Metric
CB CBs H Overall Acc. Class Acc.
Num. Loc. Per. Desc. Ent. Ab.
ZCB 14.2 16.0 18.9 77.1 80.2 92.2 88.7 69.1 64.5 74.1
ZCBs 15.0 13.6 17.7 79.0 90.0 89.3 87.7 70.0 64.5 74.1
ZH 21.3 31.7 12.6 77.3 93.8 88.5 91.3 75.6 39.7 88.9
ICL 35.9 19.9 18.5 68.5 82.9 89.7 88.7 72.9 9.9 96.3
Table 10: Comparison of TREC test set evaluation scores (%) for three DNIP objective functions ZCB, ZCBs, and ZH. Average scores over three runs of different 1-shot demonstrations are reported.
DNIP Objective (Results without DNIP are shown in the last row.) Evaluation Metric
CB CBs H Overall Acc. Class Acc.
False True
ZCB 4.5 4.5 26.0 74.1 71.8 76.3
ZCBs 4.5 4.5 26.0 74.1 71.8 76.3
ZH 13.3 13.3 24.7 75.7 68.7 82.0
ICL 43.4 43.4 29.7 71.5 48.6 92.0
Table 11: Comparison of RTE test set evaluation scores (%) for three DNIP objective functions ZCB, ZCBs, and ZH. Average scores over three runs of different 1-shot demonstrations are reported.
DNIP Objective (Results without DNIP are shown in the last row.) Evaluation Metric
CB CBs H Overall Acc. Class Acc.
Trbl. Bad Ok Gd. Grt.
ZCB 18.7 13.2 36.8 48.3 47.0 56.6 19.4 54.1 56.6
ZCBs 29.3 22.5 38.6 49.3 50.7 59.3 0.9 68.9 54.4
ZH 35.6 34.8 34.8 50.6 37.2 70.9 0.9 68.8 53.3
ICL 53.1 63.9 53.1 44.9 28.7 71.6 0.0 81.6 10.8
Table 12: Comparison of SST-5 test set evaluation scores (%) for three DNIP objective functions ZCB, ZCBs, and ZH. Average scores over three runs of different 1-shot demonstrations are reported.
DNIP Objective (Results without DNIP are shown in the last row.) Evaluation Metric
CB CBs H Overall Acc. Class Acc.
Yes No Mayb.
ZCB 41.1 31.3 38.9 63.1 74.4 61.1 12.7
ZCBs 16.1 17.1 37.1 52.4 56.9 51.5 32.7
ZH 36.3 24.3 27.8 62.7 65.9 72.0 17.6
ICL 61.2 65.0 44.0 55.1 42.2 93.7 1.8
Table 13: Comparison of PubMedQA test set evaluation scores (%) for three DNIP objective functions ZCB, ZCBs, and ZH. Average scores over three runs of different 1-shot demonstrations are reported.
Appendix C Derivation of SA Algorithm Complexity
[…]
Appendix D Additional Heatmaps for Pairwise Accuracy Differences
Figure 9 and 10 demonstrate that more balanced class accuracies are achieved with DNIP for smaller LLMs of Llama-2-7B and GPT-2-XL.
Figure 9: Comparisons of pairwise test class accuracy differences before and after applying DNIP on GPT-2-XL; the pinker the higher accuracy difference, the greener the lower accuracy difference; average accuracy over three runs are used.
Figure 10: Comparisons of pairwise test class accuracy differences before and after applying DNIP on Llama-2-7B; the pinker the higher accuracy difference, the greener the lower accuracy difference; average accuracy over three runs are used.
Appendix E Additional Ablation Analysis
We show ablation analysis on GPT-2-XL and Llama-2-7B in Figure 11 and 12.
Figure 11: Ablations on objective function, GPT-2-XL, demonstrating that $z_{1}+\beta z_{2}-\tau z_{3}$ achieves a balance point between accuracy and COBias.
Figure 12: Ablations on objective function, Llama-2-7B, demonstrating that $z_{1}+\beta z_{2}-\tau z_{3}$ achieves a balance point between accuracy and COBias.
Appendix F Full Table with Varying Number of Optimization Examples for DNIP
The full results for all three models across varying optimization set sizes are shown in Table 14.
Opt. set Eval. Metric AGNews DBpedia SST-5 TREC RTE DDI PubMedaQA
GPT-2-XL
0
Acc
COBias
52.15.4
35.511.5
31.89.9
40.03.6
34.913.7
48.75.4
27.410.5
45.68.7
55.41.9
82.424.5
14.54.4
40.75.9
55.20.0
59.412.6
10$\ast$
Acc
COBias
72.82.1
17.21.7
40.65.2
28.06.6
40.33.4
40.11.8
51.55.7
39.71.0
50.30.8
6.57.6
23.51.0
40.63.5
53.70.6
41.80.0
100
Acc
COBias
70.05.0
14.81.3
41.429.9
27.88.9
44.31.3
30.83.4
49.010.1
31.15.9
49.82.5
5.34.9
42.25.7
17.17.1
46.710.8
29.216.6
500
Acc
COBias
68.21.2
2.71.8
44.533.2
25.79.6
43.41.5
29.83.7
49.98.5
31.46.2
50.52.4
5.23.9
45.69.2
24.514.3
49.911.4
17.715.8
1,000
Acc
COBias
69.61.0
3.01.2
42.032.2
27.611.0
42.82.3
28.25.9
50.99.0
24.55.4
50.32.2
5.23.3
48.34.9
22.92.5
same as
full
Full
Acc
COBias
68.51.0
1.40.5
69.99.1
24.18.3
44.52.20
26.02.5
46.312.7
27.27.2
50.82.1
7.15.0
43.914.9
17.07.1
57.11.3
29.825.0
Llama-2-7B
0
Acc
COBias
86.42.5
14.06.5
88.92.0
13.52.1
42.111.1
55.61.5
66.76.6
33.210.0
66.34.3
61.610.5
6.70.4
41.41.7
40.36.7
40.916.1
10$\ast$
Acc
COBias
86.42.5
14.06.5
89.91.4
12.52.1
51.40.9
36.23.6
70.10.9
22.45.4
74.92.1
4.85.1
31.721.0
26.78.2
44.50.6
28.73.9
100
Acc
COBias
88.40.4
5.80.6
91.80.7
9.71.0
50.91.5
34.312.7
70.11.0
16.72.4
73.62.6
2.90.7
44.92.5
21.06.6
62.64.5
35.418.1
500
Acc
COBias
86.80.9
2.80.3
92.10.6
8.61.6
50.81.7
35.812.3
69.71.2
15.22.1
74.32.9
3.11.4
69.31.7
34.50.3
63.57.5
37.320.4
1,000
Acc
COBias
86.80.3
1.90.5
92.50.2
8.00.4
51.02.1
30.920.4
69.50.8
15.71.2
74.02.5
2.61.2
56.57.9
29.53.4
same as
full
Full
Acc
COBias
86.70.4
1.30.1
92.90.4
7.70.6
50.62.7
28.021.6
68.11.0
15.91.6
73.92.3
1.91.8
44.53.8
11.63.0
62.78.3
35.422.8
Llama-2-13B
0
Acc
COBias
79.97.0
28.316.1
88.61.7
16.23.7
44.94.3
53.15.0
68.510.8
35.96.5
71.52.2
43.47.0
7.20.9
45.65.9
55.12.9
61.21.9
10$\ast$
Acc
COBias
86.01.9/
14.33.5
88.81.6/
14.72.5
49.20.8/
41.78.1
75.83.2
29.64.5
74.82.3
12.25.7
18.712.6
30.07.8
59.66.2
35.37.8
100
Acc
COBias
87.80.2/
8.30.5
91.80.2/
10.10.7
47.42.0/
30.44.4
78.32.8
18.52.6
75.51.0
10.46.9
22.22.0
17.34.4
56.910.4
25.816.3
500
Acc
COBias
87.11.5
5.33.1
92.70.0/
9.10.3
49.50.7/
28.03.2
78.72.1
16.31.0
73.61.7
4.03.4
48.227.4
28.011.5
62.215.8
40.429.4
1,000
Acc
COBias
87.70.7/
6.71.5
92.80.1/
9.20.2
49.80.7/
28.82.0
77.32.0
14.21.8
73.91.5
3.64.3
52.010.5
25.09.7
same as
full
Full
Acc
COBias
87.90.7
6.30.6
93.40.6
7.70.6
48.31.9
18.710.1
77.12.0
14.21.3
74.30.8
4.33.3
40.46.0
7.53.2
63.114.0
41.129.6
Table 14: Comparisons of DNIP results on varying optimization set sizes; min. optimization set size is 10 for all datasets except for DBpedia, which we use 15 to cover its 14 classes. For full optimization set sizes, AGNews: 9,500, DBpedia: 9,500, SST-5: 8,116, TREC: 5,179, RTE: 2,365, DDI: 9,500, PubMedQA: 950.