### v5

# COBias and Debias: Balancing Class Accuracies for Language Models in Inference Time via Nonlinear Integer Programming

Abstract
Large language models (LLMs) are good knowledge bases but struggle to perform equally well for all classes in text classification tasks. This paper investigates a fundamental inference-time problem in language models: imbalanced class accuracies. We find what’s underneath the issue is a tendency to over-predict some classes while under-predicting some others. This class accuracy imbalance is difficult to solve from the root via better pre-training or fine-tuning strategies, but we show it can be effectively mitigated via inference-time combinatorial optimization. To this end, we conceptualize and quantify the over- and under-prediction issue as the Contextual Oddity Bias (COBias), and propose the Debiasing as Nonlinear Integer Programming (DNIP) model to correct in-context learned class probabilities based on minimizing COBias and maximizing overall accuracy, without LLM parameter update. Considering that the DNIP model implicitly contains non-differentiable elements, we therefore use the simulated annealing algorithm to solve it. Extensive evaluations on three LLMs across seven NLP classification tasks in different prompting settings show that DNIP simultaneously achieves significant COBias reduction (-27%) and accuracy improvement (+12%) over the conventional ICL approach, suggesting that inference-time mitigation of class accuracy imbalance is a promising direction to push forward LLM performances.
1 Introduction
The fact that large language models (LLMs) achieve remarkable overall test accuracy by few-shot in-context learning (ICL) [1] can obscure a large difference in individual class accuracies. Mitigating class accuracy imbalance is under-explored, but important to facilitate a smoother user experience and reduce safety risks.
It could be difficult to disentangle all the factors that cause class accuracy imbalance. Resolving the issue from the root through pre-training or fine-tuning, leveraging data selection or data augmentation, is neither easy nor cost-effective. To overcome the issue, we conceptualize a bias metric, the Contextual Oddity Bias (COBias; Section 2), to quantify class accuracy imbalance from inference-time over-predictions and under-predictions. We aim to reduce COBias by directly rectifying the in-context learned LLM output class probabilities. We propose DNIP, the Debiasing as Nonlinear Integer Programming model, to search for correction weights leveraging combinatorial optimization, without updating LLM parameters (Section 3). These weights adjust class probabilities towards lower COBias and higher overall accuracy. On three widely used different-scale LLMs across seven NLP evaluation datasets, and in different ICL settings, we show the effectiveness of DNIP.
Our key messages are as follows:
  • We introduce the COBias evaluation metric to quantify the class accuracy imbalance in language model predictions, where LLMs of different scales and families consistently exhibit a large COBias score.
  • We propose Debiasing as Nonlinear Integer Programming (DNIP) method to find class correction weights to correct ICL output class probabilities, which minimize COBias and maximize accuracy. An efficient simulated annealing algorithm is adopted to solve the nonlinear integer programming model.
  • DNIP achieves the best of both worlds of lower COBias and higher overall accuracy. For example, on Llama-2-13B, DNIP significantly reduces COBias (avg. $40.5\%\rightarrow 14.3\%$) and improves overall accuracy (avg. $59.4\%\rightarrow 69.2\%$) over the 1-shot ICL results, and it reduces COBias even further over the 5-shot ICL results (avg. $39.2\%\rightarrow 10.3\%$).
2 Over-prediction, Under-prediction, and the Contextual Oddity Bias
Pred.
World Sports Business Tech All Acc.
World 1093 64 126 3 1286 0.85
Sports 9 1247 14 0 1270 0.98
True Business 25 4 1167 8 1204 0.97
Tech 156 27 822 235 1240 0.19
All 1283 1342 2129 246 5000 -
Prec. 0.85 0.93 0.55 0.96 - -
Table 1: Over-prediction in class Business and under-prediction in class Tech in Llama-2-13B, shown by the test confusion matrix for AGNews.
Pred.
World Sports Business Tech All Acc.
World 1100 50 67 69 1286 0.86
Sports 16 1240 11 3 1270 0.98
True Business 44 2 1028 130 1204 0.85
Tech 55 7 138 1040 1240 0.84
All 1215 1299 1244 1242 5000 -
Prec. 0.91 0.95 0.83 0.84 - -
Table 2: Significantly improved test accuracy of the under-predicted class Tech for AGNews after DNIP correction.
Over-predictions and under-predictions can hurt LLM prediction capabilities. To see how, Table 2 presents an example of observed over-predictions and under-predictions in Llama-2-13B. We prompt the LLM on the AGNews news classification [2], and find that Llama-2-13B greatly over-predicts Business while under-predicting Tech, leading to low accuracy in Tech (19%). Over-/Under-prediction happens in many downstream tasks and also in larger models such as ChatGPT (Appendix A), suggesting room for improvements so that LLMs can predict all classes with equally good accuracy.
Conceptualizing Over-prediction and Under-prediction into the Contextual Oddity Bias. To introduce the Contextual Oddity Bias, we first elaborate the over-/under-prediction observations. In Table 2, the most significant contributor to the under-prediction in Tech (accuracy: 19%) is the 822 ground-truth Tech instances which are wrongly predicted as Business [^1] It simultaneously leads to a fairly low precision score of 55% in Business. In fact, the low precision score in Business and the low recall/accuracy score in Technology co-reflect each other. In other words, class accuracies are essential indicators for class performances. All current evaluation metrics are computed by Python scikit-learn. Note that by scikit-learn, class recall is the same as class accuracy.. That is, the output class Business which a ground-truth class Tech is most wrongly predicted as can reflect the model’s inference bias towards an output class for the given ground-truth class. To precisely quantify this inference bias, we first conceptualize odd classes. An odd class is an output class which holds the most wrong predictions of a ground-truth class. That is, there is always an odd class to a ground-truth class. For $N$ classes and $M$ instances, let $y$ and $\hat{y}$ denote the ground-truth class and the prediction respectively, let $u_{c_{i},c_{j}}$ denote the number of ground-truth $c_{i}$ instances that are wrongly predicted as $c_{j}$,
$u_{c_{i},c_{j}}=\displaystyle\sum\nolimits_{m}\mathbbm{1}\{y_{m}=i,\hat{y}_{m}=j\},$ (1)
$\textrm{odd}_{i}=\textrm{odd\_class}(c_{i}):=\operatorname*{arg\,max}_{c_{j}}u_{c_{i},c_{j}},$ (2)
where $m=1,...,M,i,j=1,...,N,j\neq i$, $\mathbbm{1}\{\cdot\}=1$ if the condition holds and 0 otherwise. For example, odd_class(Tech) = Business. Then, we propose to quantify how much an odd class contributes to under-predicting a ground-truth class. One may immediately use the ratio of the number of odd class predictions over that of the ground-truth instances, e.g., use 822/1240 to measure the odd Business’s “hurt” to the ground-truth Tech, but minimizing these rates alone may not adequately mitigating class accuracy imbalances (More analyses on this measurement are in Appendix B.). A simpler yet better measurement is the absolute accuracy difference between the odd class and the ground-truth class. For example, the 78% difference between Business (97%) and Tech (19%) quantifies how much Tech is under-predicted. The 19% also implicitly reflects the 822 wrong predictions in Business. It is noted that these accuracy differences sometimes fail to capture the ground-truth-and-odd relationships. For instance, World is the odd class to Business, but the former is 12% (97%-85%) below the latter. Despite the drawback, minimizing these differences will in return significantly reduce odd class predictions and improve under-predicted classes’ accuracies.
The Contextual Oddity Bias. To this end, we propose Contextual Oddity Bias (COBias). First, we propose COBiassingle, the absolute value of the accuracy difference between a ground-truth class and its odd class. Formally, given an $N$ classes and $M$ instances, let $x_{m}$ denote an instance whose label is $y_{m}$. After obtaining few-shot prompted softmax probabilities over the vocabulary at the output posistion, we take the probabilities for the class tokens and normalize them to form class probabilities, i.e., $\boldsymbol{p}_{m}=(p_{m1},\dots,p_{mN})$ is the normalized class probabilities for classes $c_{1},\dots,c_{N}$. Then the prediction $\hat{y}_{m}$ is $\operatorname*{arg\,max}_{j\in\{1,\dots,N\}}\boldsymbol{p}_{m}$. For a given class $c_{i}$, let $A_{\textrm{odd}_{i}}$ and $A_{c_{i}}$ denote the accuracies for the odd class $\textrm{odd}_{i}$ and $c_{i}$ respectively, the COBiassingle metric is computed as:
$\textrm{COBias}_{\textrm{single}}=\frac{1}{N}\displaystyle\sum\nolimits_{i}\bigg{|}A_{\textrm{odd}_{i}}-A_{c_{i}}\bigg{|},$ (3)
where class accuracy $A_{c_{i}}=\frac{1}{|\mathcal{S}_{c_{i}}|}\sum\nolimits_{m\in\mathcal{S}_{c_{i}}}\mathbbm{1}\{\hat{y}_{m}=y_{m}\}$, $\mathcal{S}_{c_{i}}$ is the set of indices for class $c_{i}$ instances. Similarly, $A_{\textrm{odd}_{i}}=\frac{1}{|\mathcal{S}_{\textrm{odd}_{i}}|}\sum\nolimits_{m\in\mathcal{S}_{\textrm{odd}_{i}}}\mathbbm{1}\{\hat{y}_{m}=y_{m}\}$, where $\mathcal{S}_{\textrm{odd}_{i}}$ is the set of indices for class $\textrm{odd}_{i}$ instances. COBiassingle measures a model’s tendency to under-predict the true answer and mostly mistake it as an odd answer. The odd classes in a same downstream task can vary across different models. That’s why we use “contextual” to mean that the oddity of an answer depends on the model or task.
Beyond the single odd classes, there could also be a secondly most-, or third most-, to the least-frequent wrongly predicted class, i.e., the odd, the second odd, to the least odd classes. For example, for ground-truth Tech, Business is the odd, while World is the second odd, and Sports is the least odd. We extend $\textrm{COBias}_{\textrm{single}}$ to every pair of classes. This arrives at the ultimate the COBias metric as follows.
$\textrm{COBias}={N\choose 2}^{-1}\displaystyle\sum_{i=1}^{N-1}\sum_{j=i+1}^{N}\bigg{|}A_{c_{i}}-A_{c_{j}}\bigg{|}.$ (4)
COBias directly reflects the inter-class imbalance in accuracy. Although the absolute value of accuracy differences of some pairs in Equation 4 may mask the odd class, it reflects a well-rounded measurement and a powerful metric when minimal class accuracy imbalance is demanded.
Figure 1: An overview of the DNIP method. We take the TREC [3] task for illustration; the ICL approach uses 1-shot prompting whose prompt is given by light yellow text boxes. The upper-left plot shows that Loc. is an odd class to the others, holding their majority wrong predictions, especially for Ent.; the upper-right plot demonstrates the stark contrast in class accuracy, with 100% for Loc. and 0 for Ent.. The bottom plots show that DNIP greatly reduces pairwise class accuracy differences, especially boosting accuracy for Ent. ($0\rightarrow 66\%$).
3 Improving LLM Outputs by Debiasing as Nonlinear Integer Programming
Essentially, we aim to rectify biased few-shot prompted LLM class probabilities with correction weights, so that corrected class probabilities should lead to less class accuracy imbalance. With this goal, we propose an inference-time nonlinear integer programming (NIP) selection model to search for optimal correction weights from a weight scale that directly minimizes COBias while maximizing overall accuracy. Therefore, the correction weights are optimized by the NIP model on a set of given LLM output class probabilities with ground-truth labels. For a test instance, after obtaining its class probabilities, the learned correction weights are applied to correct them. An overview of the proposed Debiasing as Nonlinear Integer Programming (DNIP) method is shown in Figure 1.
3.1 The DNIP Objective
We first introduce a discrete $K$-point correction weight scale $\boldsymbol{\omega}$ to select correction weights from. We denote $\boldsymbol{\omega}=(\omega_{1},\dots,\omega_{K})$, and $\omega_{k}=k/K$ is a value from the scale. For example, a $10$-point scale renders $\boldsymbol{\omega}=(0.1,\dots,1.0)$. To debias the $N$-dimensional $\boldsymbol{p}_{m}$, for each class $j$, $\boldsymbol{p}_{mj}$ multiplies a correction weight selected from $\boldsymbol{\omega}$ to obtain a corrected class probability.
Therefore, finding optimal correction weights for all $N$ classes becomes a combinatorial optimization problem: each class selects a weight from the $K$-point scale, so we need to find the optimal combination of the selections. To this end, we formulate the following correction weight selection model towards minimizing pairwise class accuracy differences and maximizing overall accuracy. Formally, we define integer selection variables $\xi=(\xi_{1},\dots,\xi_{N})$ that selects a weight index $k$ for the $n$-th class probability in $\boldsymbol{p}_{m}$, namely, $\xi_{n}=k$ if the $n$-th class selects $\omega_{k}$. As a result, the DNIP model can be given by:
$\min\quad$ $Z(\xi)=\frac{1}{M}\displaystyle\sum_{m=1}^{M}\mathbbm{1}\{\hat{y}_{m}\neq y_{m}\}+$
$\beta{N\choose 2}^{-1}\sum_{i=1}^{N-1}\sum_{j=i+1}^{N}\bigg{|}A_{c_{i}}-A_{c_{j}}\bigg{|}-\tau\displaystyle\sum_{j=1}^{N}\textrm{PMI}_{j}$ (5)
s.t. $\hat{y}_{m}=\operatorname*{arg\,max}_{n\in\{1,\dots,N\}}\omega_{\xi_{n}}p_{mn},\qquad m=1,\dots,M$ (6)
$\textrm{PMI}_{j}=\textrm{PMI}(\hat{S}^{j},S^{j})=\log\frac{f(\hat{S}^{j},S^{j})}{f(\hat{S}^{j})f(S^{j})},$
$j=1,\dots,N$ (7)
$\xi_{n}\in\{1,\dots,K\},\qquad n=1,\dots,N$ (8)
where $\beta$ and $\tau$ are parameters that control the bias and the PMI terms respectively. The PMI term is introduced to maximize the pointwise mutual information between ground-truth class $j$ instances and predicted class $j$, which can be seen as a confidence level, or a constraint to penalize each individual class. $\hat{S}^{j}$ and $S^{j}$ denote the instances of prediction $j$ and true label $j$ respectively, $f(\hat{S}^{j})$ is the ratio between the number of instances with prediction $j$ and the total number of instances, similarly, $f({S}^{j})$ is the ratio between the number of instances labeled $j$ and the total number of instances, $f(\hat{S}^{j},S^{j})$ is the ratio between the number of correct predictions of class $j$ and the total number of instances. In actual computations, we apply add-$\mu$ smoothing to the respective ratios in the numerator and denominator to smooth out zeros. The smoothing coefficient $\mu$ is selected along other parameters on a development set.
Intuitively, the first term directly minimizes error rates, the second term COBias directly minimize the inter-class accuracy differences, and the PMI term, is employed to further encourage predictions within a class to be close to the actual class. The corrected class probabilities and prediction are:
$\boldsymbol{p}^{\ast}_{m}$ $=(\omega_{\xi^{\ast}_{1}}p_{m1},\dots,\omega_{\xi^{\ast}_{N}}p_{mN})$ (9)
$\hat{y}^{\ast}_{m}$ $=\operatorname*{arg\,max}_{n\in\{1,\dots,N\}}\boldsymbol{p}^{\ast}_{m}$ (10)
3.2 Solving DNIP with Simulated Annealing
Algorithm 1 Optimizing the selection of correction weights with simulated annealing
1:$(\xi,\boldsymbol{p},y,\omega)$: weight selection variable, class probabilities and labels, the weight scale
2:$(\xi,\{y^{\ast}_{m}\}_{1}^{M})$
3:$T\leftarrow T_{max}$, $\xi^{\ast}\leftarrow\xi\leftarrow$ INIT()
4:while $T\geq T_{min}$ do
5:     while inner-loop criterion is not satisfied do
6:         $\xi_{new}\leftarrow$ PERTURB($\xi$$\triangleright$ Sample a new $\xi$
7:         $\Delta z\leftarrow$ $z(\xi_{new})-z(\xi)$
8:         if $\Delta z\leq 0$ then
9:              $\xi\leftarrow$ $\xi_{new}$
10:              if $z(\xi_{new})<z(\xi^{\ast})$ then
11:                  $\xi^{\ast}\leftarrow$ $\xi_{new}$
12:         else if random($0,1$) $<$ exp($\Delta z/T$then
13:              $\xi\leftarrow$ $\xi_{new}$triangleright Accept a worse $\xi$
14:     $T\leftarrow$ $\alpha T$
15:$y^{\ast}\leftarrow$ INFER($\xi^{\ast},\boldsymbol{p},\omega$)
16:return $\xi^{\ast},y^{\ast}$
Because the NIP objective contains discontinuous functions and is non-differentiable, we solve DNIP by simulated annealing [4, 5]. Simulated annealing (SA) is a Metropolis-Hastings sampling algorithm, which is widely adopted to solve optimization problems with its versatility in dealing with nonlinear models and many constraints [6]. The DNIP mathematical model is nonlinear, so SA is suitable. The main advantage of SA is escaping from local optima and finding global optima after reaching local optima, which is helpful in finding optimal correction weights.
In details, the SA algorithm consists of searching the solution space starting from a randomly initialized $\xi$ and generating a new one by perturbing it. The new solution is either accepted or rejected by the Metropolis criterion after evaluating the DNIP objective. Algorithm 1 describes the searching for optimal $\xi$ using SA. The outer loop performs the cooling/annealing process. In this work, the cooling schedule is set with a geometric decay with $\alpha=0.95$ and initial temperature of 200,000. The inner loop simulates the thermal equilibrium reaching process at a temperature, and the criterion ensures that either the number of generated or accepted solutions at a temperature is above a threshold [7]. We sample a new $\xi$ from the neighborhood by randomly substitute a selection in $\xi$. The computational complexity for SA is $\mathcal{O}(NK)$. Please refer to Appendix C for derivations.
4 Experiments
4.1 Experimental Setup
Evaluation Tasks. We evaluate on a diverse set of binary and multi-class NLP classification tasks, across general domains and biomedical domains. The five general-domain evaluation datasets are 4-way news topic classification AGNews, 14-way ontology classification DBpedia [8], 5-way sentiment classification SST-5 [9], 6-way retrieval question classification TREC, binary entailment recognition RTE [10]; the two domain-specific datasets from BLURB [11] are 5-way biomedical relation extraction DDI [12], and 3-way biomedical question answering PubMedQA [13]. The selected biomedical tasks are of practical uses. For example, detecting DDI is useful in preventing adverse effects from drug combinations. Each dataset comes with a training and a test set. We take the full or subset of the training set and obtain ICL class probabilities with ground-truth labels as the optimization set for DNIP optimization. Evaluations are performed on each test set against COBias and overall accuracy. Optimization and dev splits can be found in Appendix D.
Models and Experimental Configuration. We evaluate three open-source LLMs across varied scales and families, including GPT-2-XL (1.5B parameters), Llama-2-7B, and Llama-2-13B. We do few-shot prompting on an NVIDIA A100 GPU, main results use 1-shot prompting, and we also perform more few-shot analyses. For DNIP, $\beta,\tau,K,\mu$ are tuned on the development set. To account for variance caused by demonstrations in prompts, we perform three runs for each dataset with different 1-shot demonstrations (randomly selected from the optimization set). Average scores over three runs are reported. SA executes on CPU.
Model Eval. Metric AGNews DBpedia SST-5 TREC RTE DDI PubMedaQA Avg.
1-shot ICL
GPT-2-XL
Acc
COBias
52.15.4
35.511.5
31.89.9
40.03.6
34.913.7
48.75.4
27.410.5
45.68.7
55.41.9
82.424.5
14.54.4
40.75.9
55.20.0
59.412.6
38.8
50.3
Llama-2-7B
Acc
COBias
86.42.5
14.06.5
88.92.0
13.52.1
42.111.1
55.61.5
66.76.6
33.210.0
66.34.3
61.610.5
6.70.4
41.41.7
40.36.7
40.916.1
56.8
37.2
Llama-2-13B
Acc
COBias
79.97.0
28.316.1
88.61.7
16.23.7
44.94.3
53.15.0
68.510.8
35.96.5
71.52.2
43.47.0
7.20.9
45.65.9
55.12.9
61.21.9
59.4
40.5
DNIP
GPT-2-XL
Acc
COBias
68.51.0
1.40.5
69.99.1
24.18.3
44.52.20
26.02.5
46.312.7
27.27.2
50.82.1
7.15.0
43.914.9
17.07.1
57.11.3
29.825.0
54.4
18.9
Llama-2-7B
Acc
COBias
86.70.4
1.30.1
92.90.4
7.70.6
50.62.7
28.021.6
68.11.0
15.91.6
73.92.3
1.91.8
44.53.8
11.63.0
62.78.3
35.422.8
68.5
14.5
Llama-2-13B
Acc
COBias
87.90.7
6.30.6
93.40.6
7.70.6
48.31.9
18.710.1
77.12.0
14.21.3
74.30.8
4.33.3
40.46.0
7.53.2
63.114.0
41.129.6
69.2
14.3
Table 3: Performance comparisons between the 1-shot ICL approach and DNIP. In each cell, we report the average test accuracy/COBias (%) and standard deviation over three different prompts. DNIP greatly improves test accuracy and significantly lowers test COBias upon the ICL approach.
4.2 Main Results
Averaged test accuracy (in black) and COBias (in blue) with standard deviation over three runs are shown in Table 3. Seen from the ICL COBias scores, class accuracy imbalance exists in LLMs of different sizes and families. With DNIP, test COBias is significantly reduced by an absolute 29%, 23%, and 26% for the three models respectively; test accuracy is improved by 16%, 12%, and 10% respectively. For tasks whose ICL test accuracy is already in higher 80s, e.g., DBpedia, DNIP can reduce COBias while further improving the accuracy to 90s. In addition, ICL COBias is very large (82.4%) on GPT-2-XL for RTE, reflecting severe under-predictions in class False. After applying corrections, its COBias greatly reduces to 7.1%, demonstrating DNIP’s capability for balancing very imbalanced class accuracies without hurting the overall accuracy much. DNIP is also effective for biomedical tasks, bringing an average absolute reduction of 31% and 13% in COBias, and an average absolute increase of 34% and 11% in accuracy.
Dataset Time/s Variable Size
AGNews 2278 120
DBpedia 2584 420
SST-5 1499 150
TREC 1274 180
RTE 518 60
DDI 2092 150
PubMedQA 203 90
Table 4: Optimization time (seconds).
Moreover, the 13B Llama model can suffer even higher COBias than its 7B counterpart on biomedical tasks, further showing that class accuracy imbalance does not go away as model sizes increase, and DNIP is effective to balance class accuracies for LLMs of all sizes. Furthermore, we present the confusion matrix of DNIP debiased AGNews results in Table 2. Under-predictions in Tech are greatly reduced, with its accuracy significantly improved from 19% to 84%. Additionally, DNIP requires only a small optimization set of 10 examples for a fair improvement in both COBias and accuracy, and it exhibits a further, emergent COBias reduction at several thousands optimization examples.
Figure 2: Objective function and temperature curves.
We discuss performance improvements vs. the number of optimization examples needed in Appendix E.
Computational Time. For all datasets, optimization finishes in minutes. Though the optimization time varies across variable size ($N\cdot K$) and optimization set size, it finishes in dozens of minutes even for DBpedia (14 classes; Figure 2). A weight scale of 30 is good enough for most datasets evaluated.
Optimization time using weight scale 30, $\beta=2.7,\tau=0.2$ is shown in Table 4, demonstrating the practicality of DNIP as an inference-time debiasing method without incurring much costs.
4.3 More ICL Settings
Prompting Method Metric AGNews DBpedia SST-5 TREC RTE DDI PubMedQA Avg.
Acc $82.5_{2.0}$ $93.6_{1.3}$ $45.8_{6.0}$ $58.7_{23.3}$ $61.9_{16.9}$ $34.1_{42.1}$ $70.4_{7.1}$ 63.9
5-shot ICL COBias $24.4_{4.2}$ $9.0_{2.0}$ $48.0_{15.4}$ $35.4_{13.7}$ $61.3_{40.4}$ $44.4_{5.4}$ $52.2_{19.3}$ 39.2
Acc $88.5_{0.6}$ $95.8_{0.7}$ $52.9_{2.5}$ $76.6_{8.7}$ $75.8_{4.5}$ $54.1_{5.1}$ $59.9_{1.4}$ 71.9
DNIP COBias $7.0_{0.9}$ $5.7_{1.1}$ $15.8_{12.6}$ $14.4_{7.5}$ $3.1_{1.9}$ $16.5_{7.6}$ $9.6_{4.6}$ 10.3
Acc $83.5_{1.5}$ $95.2_{1.2}$ $50.3_{2.3}$ $67.0_{12.7}$ $75.0_{0.8}$ $9.7_{1.0}$ $52.3_{5.3}$ 61.9
k-shot (1 demon. per class) ICL COBias $14.9_{5.1}$ $7.0_{2.2}$ $36.3_{7.2}$ $38.2_{5.1}$ $22.5_{13.2}$ $39.7_{3.5}$ $20.9_{4.2}$ 25.6
Acc $88.7_{0.5}$ $96.6_{0.5}$ $51.3_{1.0}$ $82.7_{1.4}$ $76.7_{3.7}$ $43.6_{6.1}$ $55.3_{2.6}$ 70.7
DNIP COBias $7.3_{0.5}$ $4.3_{0.7}$ $2.8_{1.5}$ $12.1_{5.5}$ $5.0_{3.3}$ $12.5_{4.3}$ $8.7_{1.2}$ 7.5
Table 5: DNIP results in more ICL settings. Average score over three runs are reported.
We prompt Llama-2-13B in two additional settings: 5-shot, where each demonstration is randomly selected; k-shot, where k is the number of classes and each class is represented by a demonstration, randomly selected. Table 5 shows that, for example, the relative average COBias reduction by DNIP is 65% and 74% for 1-shot (Table 3) and 5-shot respectively, with relative average accuracy increase of 16% and 11%. The rationale is, demonstrations contribute to variance in outputs, but prompt engineering alone does not mitigate much the imbalance. DNIP can correct the outputs no matter what different starting points are given.
Figure 3: Test confusion matrix for DBpedia before and after applying DNIP. DNIP greatly improves ICL predictions for Nat., which is the most under-predicted class by ICL.
4.4 DNIP Significantly Balances Class Accuracies for Classification Tasks with as Many as 14 Classes
We further investigate DBpedia. By ICL, class Nat. exhibits the largest accuracy differences to others, and it is the most under-predicted class, whose odd class is Trans. and second odd class is Bldg., shown in Figure 3 (using seed 0 run). DNIP greatly reduces Trans. and Bldg. predictions for Nat.. When averaged over all three runs, test accuracy for Nat. improves from 44% to 89%, and that for Trans. increases from 78% to 87%, demonstrating DNIP’s competence in reducing class accuracy differences.
Figure 4: Pairwise class accuracy differences on test sets before and after DNIP; the pinker the higher difference, the greener the lower; averaged over three runs. DNIP achieves nearly 0 accuracy differences for many class pairs.
We also visualize pairwise class accuracy differences before and after applying DNIP to 1-shot, Llama-2-13B ICL results in Figure 4. In each heatmap, row $i$, column $j$ cell represents the absolute difference between the test accuracies of two classes in row $i,j$ , i.e., $|A_{c_{i}}-A_{c_{j}}|$. DNIP brings the accuracy differences down to close to zero for AGNews, DDI, DBpedia, and RTE. The lighter green heatmaps in the bottom row show that DNIP consistently reduces COBias across different NLP classification tasks. Furthermore, we present more heatmaps for smaller LLMs including Llama-2-7B and GPT-2-XL in Appendix F.
Figure 5: Ablation analysis of the objective function, exemplified with Llama-2-13B; test scores averaged over 7 datasets are used for the rightmost plot.
4.5 Accuracy-related Objectives vs. COBias Objective
We show that the integration of accuracy-related objectives and the COBias objective are indispensable in achieving the best of both worlds in Figure 5. DNIP objective function $z$ consists of three terms: the error rate term $z_{1}=\frac{1}{M}\sum_{m=1}^{M}\mathbbm{1}\{\hat{y}_{m}\neq y_{m}\}$; the COBias term $z_{2}={N\choose 2}^{-1}\sum_{i=1}^{N-1}\sum_{j=i+1}^{N}|A_{c_{i}}-A_{c_{j}}|$; and the PMI term $z_{3}=\sum_{j=1}^{N}\textrm{PMI}_{j}$; the accuracy-related terms are $z_{1}$ and $z_{3}$. We ablate all ablation objective functions to perform DNIP: $z_{1},z_{2},z_{3},z_{1}+\beta z_{2},z_{1}-\tau z_{3},\beta z_{2}-\tau z_{3},z_{1}+\beta z_{2}-\tau z_{3}$. Ablations on GPT-2-XL and Llama-2-7B show similar findings (Appendix G). To make fair comparisons, we fix $\beta=2.7,\tau=0.2$ and weight scale $K=30$, which may not result in the same scores reported in Table 3. Among all ablations, $z_{1}+\beta z_{2}-\tau z_{3}$ achieves a balance point between accuracy and COBias. For the criteria, lower COBias is our top priority; when COBias scores are similarly low, we prefer the one that achieves higher accuracy. Compared to $z_{1}+\beta z_{2}-\tau z_{3}$$z_{1}$: (1) Using $z_{1}$ or $z_{3}$ solely emphasizes improving accuracy; though obtaining higher or similar accuracy, they suffer a higher COBias. (2) Using $z_{1}-\tau z_{3}$ also obtains a higher/similar accuracy but a higher COBias. (3) Using $z_{2}$ solely focuses on mitigating COBias and is straightforwardly effective in reducing COBias, but the shortcoming is a lower accuracy. (4) Using $z_{1}+\beta z_{2}$ or $\beta z_{2}-\tau z_{3}$ is promising with a similar/lower accuracy and a lower COBias. Accuracy-related objectives alone increase accuracy by over 16% ($75\%-59\%$), suggesting DNIP is effective for those who emphasize accuracy improvement.
Figure 6: Comparisons to more methods.
4.6 Optimizing COBiassingle
Using COBiassingle (Eq. 3) for DNIP optimization will focus on odd classes only. Figure 7 shows ablation results when replacing the COBias term in the DNIP objective with COBiassingle. Blue dots represent the respective evaluation metric score of the original 1-shot ICL Llama-2-13B outputs of the seven benchmark datasets; green dots are debiased results optimized by the COBiassingle DNIP objective; red dots are results optimized by the COBias DNIP objective.
Figure 7: Test COBiassingle, COBias, and overall accuracy scores (from left to right).
Comparing DNIP with COBiassingle and DNIP with COBias, their performances are close, but DNIP with COBias achieves higher overall test accuracy and slightly better bias mitigation performances. In other words, DNIP with COBiassingle can significantly minimize accuracy differences between odd classes and ground-truth classes, but DNIP with COBias is even more effective when it comes to balancing all pairs of classes, not limited to the odd classes.
5 Related Work
Class accuracy imbalance is one of the potential failures of LLMs. Analysis towards understanding LLM failure modes has gained remarkable insights. For example, understanding hallucinations in LLMs facilitates fact-checking metrics for LLMs.[14, 15, 16]. Jones and Steinhardt [17] reveals failure modes of LLMs which resemble cognitive biases. Lu et al. [18] identify order sensitivity in prompting. The class accuracy imbalance has multi-faceted causes, from pre-training data, such as predilections for common tokens in pretraining [19], to a liking for specific patterns or answers [20]. Other causes include choices of prompt templates [21, 22, 23], demonstrations [24], and prompt orders [25]. Unintended LLM behaviors may also exacerbate it, such as sycophancy[26, 27, 28].
6 Discussion
More Methods That Potentially Improve Accuracy while Reducing COBias. Prior methods tackle the prediction bias problem from different angles. Although COBias is not be explicitly modeled, they are promising to reduce COBias. Therefore, we draw comparisons to two popular categories of debiasing methods, adaptations and calibrations. In particular, with the same class probability vectors as input, we build a single linear adapter (LA) network with adapter dimension of 64, and train it towards the same objective as DNIP; we also follow the batch calibration method [29] to estimate a calibration term using test examples. Figure 6 shows the comparison results. fine-tuning a simple linear adapter does not work well on the class output probabilities. The proposed method outperforms both methods in accuracy and COBias.
Evaluating Closed LLMs. While closed LLMs such as ChatGPT [30] are interesting to debias, ChatGPT returns log probabilities of top 20 most likely tokens, which may not cover the probability for every class that is needed as inputs to DNIP.
Appendix H has more discussions on optimization algorithms for integer programming, how our motivation is different from other classical post-hoc correction methods, using calibration methods to mitigate ICL biases, and using Chain-of-Thought prompting to improve LLM predictions.
7 Conclusion
We introduce COBias to quantify class accuracy imbalance for language models, and emphasize the need to minimize COBias to push for more reliable predictions while maintaining good overall accuracy. We correct ICL output class probabilities towards this goal as a combinatorial optimization problem, proposing DNIP to optimize class correction weights. Evaluations on three LLMs across seven classification tasks demonstrate that DNIP achieves both lower COBias and higher overall accuracy over original ICL results.
Appendix A Other Examples of Over-prediction and Under-prediction in LLMs
We present Llama-2-13B’s and ChatGPT’s over- and under-predictions on a class imbalanced classification task, the Drug-Drug Interaction (DDI) relation classification task, where around 90% test instances are Negative. As shown in Table 6, Llama-2-13B significantly under-predicts Negative. In Table 7, similar under-predictions in the DDI Negative class happens for ChatGPT (GPT-3.5-turbo-instruct [31]), the RLHF’d chat LLM with magnitudes of more parameters. (It is noted that at the time of our experiments, ChatGPT does not support returning full probabilities over the entire vocabulary. To investigate the over-/under-prediction issue, we prompt ChatGPT on the DDI task with the following prompt template: “Please choose one most suitable answer for the following description. [ONE-SHOT DEMONSTRATION] Description: [TEST INPUT] Q: What is the drug-drug interaction relation between the aforementioned @DRUG\$ pair? Choose from Negative, Mechanism, Effect, Advice, Interaction. A: ”, and use the answer token directly as the prediction, and plot the confusion matrix in Table 7. For fair comparisons, we prompt Llama-2-13B in the same manner to obtain Table 6, which is different from how we prompt Llama-2-13B in the main paper.)
Pred.
Neg Eff Mech Adv Int
Neg 1749 438 1498 546 506
Eff 118 33 110 52 47
True Mech 77 28 123 45 29
Adv 84 15 62 36 24
Int 34 11 28 13 10
Table 6: Over-prediction and under-prediction in Llama-2-13B, as shown by the confusion matrix on DDI test set.
Pred.
Neg Eff Mech Adv Int
Neg 809 707 1001 217 2003
Eff 45 109 79 6 121
True Mech 15 46 172 0 69
Adv 24 12 18 79 88
Int 9 9 7 1 70
Table 7: Over-prediction and under-prediction in ChatGPT, as shown by the confusion matrix on DDI test set.
The over-/under-predictions in LLMs of different sizes and families, and the inconveniences in training/fine-tuning LLMs, together call for a more actionable bias metric quantifying the over-/under-prediction issue and an accompanying debiasing method tailored for LLMs.
Appendix B Directly Optimizing the Odd Prediction Rate
We discuss another option of measuring class accuracy imbalances, which does not mitigate odd predictions as effective as using COBias, though using it can obtain similar or slightly higher overall accuracy. It is the odd prediction rate, i.e., the rate of odd predictions over the number of ground-truth instances, which is a measurement reflecting the odd class and the level of “accuracy hurt” caused by the odd class. We define $H(c_{i})$ as below.
$H(c_{i})=\frac{\displaystyle\sum\nolimits_{m}\mathbbm{1}\{y_{m}=c_{i},\hat{y}_{m}=\textrm{odd}_{i}\}}{\displaystyle\sum\nolimits_{k}\mathbbm{1}\{y_{k}=c_{i}\}},$ (11)
where $\textrm{odd}_{i}$ is the odd class to $c_{i}$. For example, one may quantify the effect of a current odd class Business to the ground-truth class Tech as the rate of Business predictions over ground-truth Tech instances, i.e., 822/1240, and an immediate debiasing goal is to minimize this rate. The DNIP objective with the odd prediction rate is as follows.
$\min\quad Z(\xi)=\frac{1}{M}\displaystyle\sum_{m=1}^{M}\mathbbm{1}\{\hat{y}_{m}\neq y_{m}\}+\beta\frac{1}{N}\displaystyle\sum\nolimits_{i}H(c_{i})-\tau\displaystyle\sum_{j=1}^{N}\textrm{PMI}_{j}$ (12)
We compare the effectiveness of using COBias, COBiassingle, and $H$ in the DNIP objective, and refer to the respective objectives as ZCB (Eq. 3), ZCBs, and ZH (Eq. 12). Detailed test scores for the benchmark datasets on Llama-2-13B, including class accuracies, are shown in Tables 8 to 14.
In short, ZH still leaves some classes under-predicted, as examplified by the individual test class accuracies, and test COBias (or test COBiassingle) scores on TREC, RTE, etc. The rationale is that optimizing bias measurements that extend accuracy differences, such as the odd prediction rates, does not mitigate odd class predictions better than optimizing COBias directly. ZCB most effectively mitigates the over- and under-prediction issue, balancing class accuracies.
Optimizing other extended measurements, such as precision or $F_{1}$ scores differences, may obtain similar performances to the three existing metrics, which will be left for future discussions.
DNIP Objective (Results without DNIP are shown on the last row.) Evaluation Metric
CB CBs H Overall Acc. Class Acc.
Co. Sch. Art. Ath. Polit. Trans. Bldg. Nat. Vill. Anim. Plt. Alb. Film Bk.
ZCB 7.7 7.5 3.3 93.4 82.6 97.7 94.0 99.5 97.9 94.4 86.6 89.0 99.5 96.3 98.5 98.6 97.0 77.8
ZCBs 9.1 5.7 3.9 92.6 87.4 96.8 97.1 98.9 98.1 87.3 87.6 84.0 99.3 99.2 97.9 99.6 96.5 69.0
ZH 8.9 9.5 2.8 93.2 76.2 99.3 95.0 99.4 98.8 96.9 82.8 89.8 99.5 98.4 98.9 99.8 96.1 76.3
ICL 16.2 13.7 5.6 88.6 91.7 99.1 92.7 99.7 97.0 94.8 78.4 44.2 99.7 96.9 99.3 99.8 95.2 63.3
Table 8: Comparison of DBpedia test set evaluation scores (%) for three DNIP objective functions ZCB, ZCBs, and ZH. Average scores over three runs of different 1-shot demonstrations are reported.
DNIP Objective (Results without DNIP are shown on the last row.) Evaluation Metric
CB CBs H Overall Acc. Class Acc.
Wrd. Sports Bus. Tech.
ZCB 6.3 4.4 7.2 87.9 85.3 95.9 86.5 83.9
ZCBs 7.3 4.3 6.9 88.4 84.8 98.2 85.8 84.7
ZH 8.5 4.8 6.5 88.3 86.8 98.7 84.2 83.2
ICL 28.3 22.5 14.6 79.9 81.5 98.7 91.0 48.0
Table 9: Comparison of AGNews test set evaluation scores (%) for three DNIP objective functions ZCB, ZCBs, and ZH. Average scores over three runs of different 1-shot demonstrations are reported.
DNIP Objective (Results without DNIP are shown in the last row.) Evaluation Metric
CB CBs H Overall Acc. Class Acc.
Neg. Eff. Mech. Adv. Int.
ZCB 7.5 7.0 39.6 40.4 42.1 31.6 32.8 33.9 31.6
ZCBs 19.9 16.8 42.1 22.9 19.9 34.2 37.3 39.8 44.8
ZH 43.6 44.6 30.9 21.7 17.6 12.6 47.2 85.5 33.3
ICL 45.6 71.8 59.4 7.2 0.0 30.7 61.5 24.1 43.5
Table 10: Comparison of DDI test set evaluation scores (%) for three DNIP objective functions ZCB, ZCBs, and ZH. Average scores over three runs of different 1-shot demonstrations are reported.
DNIP Objective (Results without DNIP are shown in the last row.) Evaluation Metric
CB CBs H Overall Acc. Class Acc.
Num. Loc. Per. Desc. Ent. Ab.
ZCB 14.2 16.0 18.9 77.1 80.2 92.2 88.7 69.1 64.5 74.1
ZCBs 15.0 13.6 17.7 79.0 90.0 89.3 87.7 70.0 64.5 74.1
ZH 21.3 31.7 12.6 77.3 93.8 88.5 91.3 75.6 39.7 88.9
ICL 35.9 19.9 18.5 68.5 82.9 89.7 88.7 72.9 9.9 96.3
Table 11: Comparison of TREC test set evaluation scores (%) for three DNIP objective functions ZCB, ZCBs, and ZH. Average scores over three runs of different 1-shot demonstrations are reported.
DNIP Objective (Results without DNIP are shown in the last row.) Evaluation Metric
CB CBs H Overall Acc. Class Acc.
False True
ZCB 4.5 4.5 26.0 74.1 71.8 76.3
ZCBs 4.5 4.5 26.0 74.1 71.8 76.3
ZH 13.3 13.3 24.7 75.7 68.7 82.0
ICL 43.4 43.4 29.7 71.5 48.6 92.0
Table 12: Comparison of RTE test set evaluation scores (%) for three DNIP objective functions ZCB, ZCBs, and ZH. Average scores over three runs of different 1-shot demonstrations are reported.
DNIP Objective (Results without DNIP are shown in the last row.) Evaluation Metric
CB CBs H Overall Acc. Class Acc.
Trbl. Bad Ok Gd. Grt.
ZCB 18.7 13.2 36.8 48.3 47.0 56.6 19.4 54.1 56.6
ZCBs 29.3 22.5 38.6 49.3 50.7 59.3 0.9 68.9 54.4
ZH 35.6 34.8 34.8 50.6 37.2 70.9 0.9 68.8 53.3
ICL 53.1 63.9 53.1 44.9 28.7 71.6 0.0 81.6 10.8
Table 13: Comparison of SST-5 test set evaluation scores (%) for three DNIP objective functions ZCB, ZCBs, and ZH. Average scores over three runs of different 1-shot demonstrations are reported.
DNIP Objective (Results without DNIP are shown in the last row.) Evaluation Metric
CB CBs H Overall Acc. Class Acc.
Yes No Mayb.
ZCB 41.1 31.3 38.9 63.1 74.4 61.1 12.7
ZCBs 16.1 17.1 37.1 52.4 56.9 51.5 32.7
ZH 36.3 24.3 27.8 62.7 65.9 72.0 17.6
ICL 61.2 65.0 44.0 55.1 42.2 93.7 1.8
Table 14: Comparison of PubMedQA test set evaluation scores (%) for three DNIP objective functions ZCB, ZCBs, and ZH. Average scores over three runs of different 1-shot demonstrations are reported.
Appendix C Derivation of SA Algorithm Complexity
[…]
In addition, the brute force strategy to find optimal correction weights is enumerated search, where the number of weight combinations to explore is $K^{N}$ for an example; for $M$ examples, it results in a computational complexity of $\mathcal{O}(K^{N}\cdot M)$, which can not finish in polynomial time. When $M$ or $N$ is large, enumerated search is computationally intensive. Considering this, we did not adopt the brute force strategy.
Appendix D Dataset Preprocessing
We take the full training set and obtain ICL class probabilities with ground-truth labels for DNIP optimization, except for AGNews, DBpedia, and DDI, for which we randomly select 10,000 training examples, since the datasets are quite large. We further split the training set into 95% optimization set and 5% development set. Test evaluations are performed on each test set against overall accuracy and COBias, except for AGNews, for which we evaluate on 5,000 randomly selected test examples.
Appendix E Performance Improvements vs. Number of Optimization Examples
We evaluate DNIP with varying optimization examples on Llama-2-13B. We find that DNIP requires only a small optimization set of 10 examples for a fair improvement in both COBias and accuracy, and it exhibits a further, emergent COBias reduction at several thousands optimization examples.
Figure 8: Relative test COBias reduction; higher is better.
Test COBias reduction reaches 9% to 72% at 10 optimization examples, and continues to grow with more optimization examples. Figure 8 shows the relative test COBias reduction over ICL results. At 10 examples, DNIP can achieve a relative COBias reduction ranging from 9% to 72%. We observe an overall trend of COBias reduction except for DDI and PubMedQA, who have a spike at 100 examples, but DDI rises again after 500 examples while PubMedQA levels off (PubMedQA full optimization size: 950). An explanation is that DNIP prioritizes COBias reduction more than accuracy increase at fewer examples; at more examples, the accuracy-related objectives show more influence. It also indicates that domain-specific datasets are more sensitive to the trade-off. We further notice a surge in reductions from 1,000 to full for DDI, SST-5, and DBpedia (full optimization sizes: 9,500, 8,116, 9,500), suggesting that around 8,000 examples are needed for a further emergent reduction - PubMedQA may see further COBias reduction if more optimization examples are available.
Test accuracy is improved with only 10 optimization examples, and plateaus at around 500 examples. Figure 9 shows that, at 10 examples, the relative increase in accuracy ranges from 0.2% to 11% except DDI, which is 160%. At 100 examples, the increase ranges from 3% to 208% (DDI, $7.2\%\rightarrow 18.7\%$). The increase continues to grow and stabilizes at around 500 examples. At 1,000 examples, test accuracy for DDI rises to 52%, which is more than 6 times of the ICL accuracy, showing that DNIP is effective for this challenging biomedical NLP task.
Figure 9: Relative test accuracy increase; higher is better.
Moreover, we observe for RTE a declining trend of accuracy increase, which matches its growing trend of COBias reduction, reflecting the trade-off between the two objectives. We also note a slight decrease in accuracy improvement for SST-5 and DDI at full size, which matches their sudden increase in COBias reduction, verifying that DNIP has emergent COBias reduction abilities with thousands of optimization examples. Overall, DNIP improves both accuracy and COBias at 10 examples, and achieves acceptable COBias reduction and accuracy improvement at 100 examples, with test COBias reduced by 57% and test accuracy improved. For detailed scores, please refer to Appendix I.
Appendix F Additional Heatmaps for Pairwise Accuracy Differences
Figure 10 and 11 demonstrate that more balanced class accuracies are achieved with DNIP for smaller LLMs of Llama-2-7B and GPT-2-XL.
Figure 10: Comparisons of pairwise test class accuracy differences before and after applying DNIP on GPT-2-XL; the pinker the higher accuracy difference, the greener the lower accuracy difference; average accuracy over three runs are used.
Figure 11: Comparisons of pairwise test class accuracy differences before and after applying DNIP on Llama-2-7B; the pinker the higher accuracy difference, the greener the lower accuracy difference; average accuracy over three runs are used.
Appendix G Additional Ablation Analysis
We show ablation analysis on GPT-2-XL and Llama-2-7B in Figure 12 and 13.
Figure 12: Ablations on objective function, GPT-2-XL, demonstrating that $z_{1}+\beta z_{2}-\tau z_{3}$ achieves a balance point between accuracy and COBias.
Figure 13: Ablations on objective function, Llama-2-7B, demonstrating that $z_{1}+\beta z_{2}-\tau z_{3}$ achieves a balance point between accuracy and COBias.
Appendix H More Discussions
Optimization Algorithms for Integer Programming. The classic solvers for integer programming include operational research techniques such as Branch and Bound (BnB), commonly used for linear integer programming problems. However, it is difficult for these methods to solve our DNIP model, which contains discontinuous functions and is non-differentiable. In this case, a series of metaheuristic algorithms are more suitable, solversincluding Simulated Annealing (SA), Genetic Algorithms, Ant Colony Optimization, and Particle Swarm Optimization. These algorithms all belong to the same category, and each meta heuristic has their own pros and cons. Therefore, SA in this paper has a clear purpose of tackling our NIP mathematical model that classical operational research methods cannot solve.
Our Motivation Is Different from Classical Post-hoc Corrections. Some may contend that achieving equitable accuracies across all classes is a well-explored issue in standard machine learning classifiers. However, it is crucial to see that class prediction accuracy imbalances must be understood within their specific context. The accuracy bias in LLMs’ outputs arises from fundamentally different causes than the unequal class accuracies seen in potentially overfitted traditional classifiers. In the former, biases stem from prompts and pre-training, while in the latter, the imbalance is driven by skewness in the supervised training data. This distinction underscores why our approach is uniquely applied to LLMs’ output token class probabilities.
Using Calibration Methods to Mitigate ICL Biases. Previous calibration methods reduce biases caused by prompting. For example, Zhao et al. [19], Zhou et al. [29] use content-free or content-based prompts to measure an offset and calibrate class probabilities by removing it. This may be good for mitigating bias towards a single class, but it does not consider the interaction between class predictions, e.g., pairwise class accuracy bias studied in this work. Another hazard is using prompts to compute the offset, which may cascade biases brought by a certain or several prompt templates, and biases inherited from pretraining may not be reflected. Instead, we correcct biased predictions without more prompting, but we differ in that our correction weights are explicitly optimized on COBias and accuracy metrics, while the cluster-based decision boundary is implicitly learned.
Using Chain-of-Thought Prompting to Improve LLM Predictions. Some may argue that a simple remedy for biased predictions is Chain-of-Thought (CoT) prompting, which elicit more reasoning in LLMs. However, Turpin et al. [25] points out that CoT explanations can be unfaithful, as they can be biased towards class A when prompts always put letter A as the answer, causing 36% accuracy drops on BBH [32]. We add that prompt-based models may fall for pattern matching instead of recalling learned factual knowledge [20]. Therefore, CoT may not help due to unfaithful explanations.
Appendix I Full Table with Varying Number of Optimization Examples for DNIP
As an addition to Appendix E, the full results for all three models across varying optimization set sizes are shown in Table 15.
Opt. set Eval. Metric AGNews DBpedia SST-5 TREC RTE DDI PubMedaQA
GPT-2-XL
0
Acc
COBias
52.15.4
35.511.5
31.89.9
40.03.6
34.913.7
48.75.4
27.410.5
45.68.7
55.41.9
82.424.5
14.54.4
40.75.9
55.20.0
59.412.6
10$\ast$
Acc
COBias
72.82.1
17.21.7
40.65.2
28.06.6
40.33.4
40.11.8
51.55.7
39.71.0
50.30.8
6.57.6
23.51.0
40.63.5
53.70.6
41.80.0
100
Acc
COBias
70.05.0
14.81.3
41.429.9
27.88.9
44.31.3
30.83.4
49.010.1
31.15.9
49.82.5
5.34.9
42.25.7
17.17.1
46.710.8
29.216.6
500
Acc
COBias
68.21.2
2.71.8
44.533.2
25.79.6
43.41.5
29.83.7
49.98.5
31.46.2
50.52.4
5.23.9
45.69.2
24.514.3
49.911.4
17.715.8
1,000
Acc
COBias
69.61.0
3.01.2
42.032.2
27.611.0
42.82.3
28.25.9
50.99.0
24.55.4
50.32.2
5.23.3
48.34.9
22.92.5
same as
full
Full
Acc
COBias
68.51.0
1.40.5
69.99.1
24.18.3
44.52.20
26.02.5
46.312.7
27.27.2
50.82.1
7.15.0
43.914.9
17.07.1
57.11.3
29.825.0
Llama-2-7B
0
Acc
COBias
86.42.5
14.06.5
88.92.0
13.52.1
42.111.1
55.61.5
66.76.6
33.210.0
66.34.3
61.610.5
6.70.4
41.41.7
40.36.7
40.916.1
10$\ast$
Acc
COBias
86.42.5
14.06.5
89.91.4
12.52.1
51.40.9
36.23.6
70.10.9
22.45.4
74.92.1
4.85.1
31.721.0
26.78.2
44.50.6
28.73.9
100
Acc
COBias
88.40.4
5.80.6
91.80.7
9.71.0
50.91.5
34.312.7
70.11.0
16.72.4
73.62.6
2.90.7
44.92.5
21.06.6
62.64.5
35.418.1
500
Acc
COBias
86.80.9
2.80.3
92.10.6
8.61.6
50.81.7
35.812.3
69.71.2
15.22.1
74.32.9
3.11.4
69.31.7
34.50.3
63.57.5
37.320.4
1,000
Acc
COBias
86.80.3
1.90.5
92.50.2
8.00.4
51.02.1
30.920.4
69.50.8
15.71.2
74.02.5
2.61.2
56.57.9
29.53.4
same as
full
Full
Acc
COBias
86.70.4
1.30.1
92.90.4
7.70.6
50.62.7
28.021.6
68.11.0
15.91.6
73.92.3
1.91.8
44.53.8
11.63.0
62.78.3
35.422.8
Llama-2-13B
0
Acc
COBias
79.97.0
28.316.1
88.61.7
16.23.7
44.94.3
53.15.0
68.510.8
35.96.5
71.52.2
43.47.0
7.20.9
45.65.9
55.12.9
61.21.9
10$\ast$
Acc
COBias
86.01.9/
14.33.5
88.81.6/
14.72.5
49.20.8/
41.78.1
75.83.2
29.64.5
74.82.3
12.25.7
18.712.6
30.07.8
59.66.2
35.37.8
100
Acc
COBias
87.80.2/
8.30.5
91.80.2/
10.10.7
47.42.0/
30.44.4
78.32.8
18.52.6
75.51.0
10.46.9
22.22.0
17.34.4
56.910.4
25.816.3
500
Acc
COBias
87.11.5
5.33.1
92.70.0/
9.10.3
49.50.7/
28.03.2
78.72.1
16.31.0
73.61.7
4.03.4
48.227.4
28.011.5
62.215.8
40.429.4
1,000
Acc
COBias
87.70.7/
6.71.5
92.80.1/
9.20.2
49.80.7/
28.82.0
77.32.0
14.21.8
73.91.5
3.64.3
52.010.5
25.09.7
same as
full
Full
Acc
COBias
87.90.7
6.30.6
93.40.6
7.70.6
48.31.9
18.710.1
77.12.0
14.21.3
74.30.8
4.33.3
40.46.0
7.53.2
63.114.0
41.129.6
Table 15: Comparisons of DNIP results on varying optimization set sizes; min. optimization set size is 10 for all datasets except for DBpedia, which we use 15 to cover its 14 classes. For full optimization set sizes, AGNews: 9,500, DBpedia: 9,500, SST-5: 8,116, TREC: 5,179, RTE: 2,365, DDI: 9,500, PubMedQA: 950.